A Complete Bibliography of ACM Transactions on Autonomous and Adaptive Systems (TAAS)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 October 2025 Version 1.46

Title word cross-reference

k [PPCE22]. p [BDMP12]. Q [ZLHC21].

* [KGL21].

-Coverage [PPCE22]. -persistent [BDMP12]. -values [ZLHC21].

2012 [Edi14]. 2013 [PH15]. 2014 [BE16]. 2015 [SI17]. 2016 [CPS17]. 2020 [TWS20]. 2021 [PK23]. 2022 [DGB24, SCM24]. 2023 [MCJ25].

3.0 [ADAR25, QKSV25, YZS^{+}25]. 3.0-Enabled [YZS^{+}25]. 3PC [HSM^{+}12].

61499 [FSF24].

802.15.4 [BRDA16].

Access [VMG14, KM08]. Accesses [WVT^{+}17]. Accurate [JH13]. Achieving [HL13]. ACM [Bah24, FMVC14, Bah24]. acquisition [FFJ^{+}12]. ACSOS [TWS20, DGB24, PK23]. Action [BSK20, Gab11, RW20, TSLF23]. Activity [CLRC21]. Actuation [MM17]. actuator [MPBMP^{+}10]. Ad [CW11, MR11, SZB19, PRJ11, SLJS08]. Adaptare [DCL^{+}12]. Adaptation [AMG18, BVPD17, Buc19, CMGS16, CLSS^{+}13, EYCM16, FMVC14, GGPTRC16, GSB24, GW24, GCA^{+}24, HBMA21, KKK^{+}16, LHKK23, LZL^{+}24b, MCGS18, RMKM17, WGQV22, WGA^{+}23, ZNU21, CMS25, DCL^{+}12, KGJ12, PSB^{+}12, RTH19, ZP12, TSLF23]. Adaptations [PBW^{+}20]. adaptative [HKR08]. Adapting [HAMR13, RTN^{+}17, CSG^{+}24]. Adaptive [AA16, APSM18, ARS17, Bah24, Bak11, BSS^{+}14, BUL^{+}18, BWW^{+}17, BRDA16, BCC^{+}23, CHC16, CY07, CLRC21, DBDF19, FSF24, FMA^{+}17, GW24, GEB^{+}21, HSL^{+}07, DW15, IJDZ16, KCH14, LZ13, LVP15, LEC^{+}15, LZL^{+}24a, LCT^{+}18, LXX^{+}14, MVV14, MLFP24, NCKA25, ONC17, PLQ^{+}25, PPCE22, PAS^{+}25, SMK21, SGP13, ST13, SHRB13, SQX^{+}07, SPB24, SKA20, VMG14, VG14, WCW^{+}17, WHA^{+}25, WLC^{+}25, XLXZ14, ZHSP20, ZSLG16, AGLV10, BDMP12, BN12, FRL09, GPTW13, GWQ21, HSM^{+}12, HWA^{+}20, KGJ08, LPZZ09, MIRG06, PDCE11, ST09, SWM19, WXZ10, WMA12, WGQV22, XVYH11, YTW08, ZS09]. Adaptiveness [PSPR15]. Adaptivity [LNS24, SMSÇ^{+}10]. Added [AMS^{+}19]. Adding [CEK14]. Admission [GSD08]. Advantage [BMS11]. Aerial [CHCZ^{+}24, PAS^{+}25]. affect [KIW06]. against [DXP14, KD07, LXX^{+}14]. Agent [AA16, ADV16, BRII21, CLSS^{+}13, CW14, EGSM21, GR08, LCQB19, LV07, MDC17, PPB17, WCW^{+}17, XGB^{+}25, DDH23, FRL09, GCC06, GZBC25, HKR08, KGJ12, MALS22, Pos07, WHH^{+}10b, WHH10a, ZLHC21]. Agent-based [EGSM21, GCC06]. Agentic [PCL^{+}25]. AgentLink [POPM07]. Agents [BEE^{+}20, Buc19, DBDF19, FCD^{+}18, RH16, SZB19, SZY^{+}20, JI07, MIRG06, PPSM07]. Aggregates [FHDD23]. Agile [USC^{+}08]. Agilla [FRL09]. Ahead [ZLS22]. AI [FPC24, LZL^{+}24a, LHAES23, PCL^{+}25]. AID [ZS09]. Airborne [BUL^{+}18]. Algorithm [BMPB23, BRDA16, RPG^{+}15, SBMM17, APA12, MHZ13, SLJS08]. Algorithmic [WV18]. Algorithms [BSS^{+}14, CFGM16, DP16, FMSA11, Gab11, SCC17, BN12, DKMD11, FMS08, KGJ08, PB13, WDTS11]. Alleviation [PLQ^{+}25]. Allocation [ASS^{+}15, AH21, BRII21, FLHP24, JZL15, KPO19, ZCS12]. Ambient [CLRC21, AGLV10, DHC10, Her10, LV10]. A}nalysis [BCF^{+}08, ALKG22, GGPTRC16, KTK^{+}16, SGP13, SBMM17, Lit07]. Analytics [LF19]. Analyzing [CMGS16, WNV12a]. Android [RMM19]. Anomaly [FRA^{+}25, IJDZ16, ZS09]. Anonymity [BR11]. Anonymous [DK12, MT09]. Ant [SLJS08, LDD06, PB13]. Ant-based [SLJS08]. Anticipatory [ZHSP20]. Anunnaki [LZC24]. Anycast [CW11]. Application [GCC06, LR12, PSB^{+}12]. Applications [BDLM11, FBL18, GB14, HBDD14, LCT^{+}18, LDC^{+}18, MS15, MVV14, QKSV25, SHRB13, WXG^{+}25, ZSLG16, AGLV10, CSLZ10, LS09, USC^{+}08]. Applying [BNL24, GWQ21]. Approach [CMRZ15, CLW^{+}14, DLPT14, EK21, Gab11, JAJ^{+}18, LC21, MVV14, Men16, PPCE22, SWM19, SZB19, VMG14, VGR^{+}15, WVT^{+}17, AVC09, GPTW13, GYSD08, GR10, LV07, PSB^{+}12, PSFC12, XVYH11]. Approaches [FRA^{+}25, KGJ12]. Approximation [NCKA25, PB13]. Apps [RMM19]. Arbitrary [MG11]. Architectural [WHA^{+}25]. Architecture [FDMD15, HKR08]. Architectures [GGV20]. Area [HBDD14]. Argumentation [SDOP20]. Argumentation-Based [SDOP20]. Array [PSPR15]. Arrays [HMF^{+}15]. Art [LZL^{+}24a]. Articles [FMVC14]. Artificial [CHC16, LZC24, WHY24]. Assembly [FP17, GR10, TGT^{+}06]. Assignment [RW20]. association [DHC10]. Assured [ZNU21]. Asynchrony [BR11]. AT-DIFC [SKA20]. Attack [TOLG^{+}24]. Attacks [BSS^{+}14]. Attributed [HEC^{+}16]. Auction [TVKB16, ZCS12]. Augmented [IJDZ16]. Authentication [BCC^{+}23]. Auto [DRPQ14, LDC^{+}18]. Auto-Profiling [LDC^{+}18]. AutoHome [BDLM11]. Automated [FE12, KIW06, MLsRA^{+}15, PLQ^{+}25, DDH23]. Automatic [SC19, CLHX12, DCL^{+}12]. automatically [YTW08]. Automaton [MG11]. Autonomic [AOK11, BBC^{+}11, BDLM11, DXP14, DLPT14, DC12, IA18, LZ13, LF19, MHP^{+}12, TWS20, AVC09, DDF^{+}06, KM08, LS09, Lit07, XVYH11]. Autonomous [Bah24, BCC^{+}17, BWW^{+}17, BYH^{+}25, CHZ^{+}25, DAHS21, FYC^{+}24, FCD^{+}18, Gab11, LTS^{+}25, MPBMP^{+}10, QKSV25, SMHP15, WLC^{+}25, XGB^{+}25, WNV12a, WNV12b]. AutoPlacer [PRRR15]. Availability [CLRC21, PLQ^{+}25, RDKB15, LLL12]. Aware [CMGS16, FCD^{+}18, MCGS18, SKA20, WHA^{+}25, DZJ^{+}21, DBA^{+}21, FS09, GSD08, LHKK23, PBW^{+}20, RTH19, WXG^{+}25]. Awareness [LV07, ESBT19, PRJ11]. Axiomatization [PSA12].

backend [SA06]. Balancing [BGW24, JZL15, AHM09, GYP12]. base [LR12]. Based [AA16, BBDB15, CLW^{+}14, CMP13, GSB24, IA18, KKK^{+}18b, LZ13, LTS^{+}25, MS15, MM17, MDC17, MBB11, SDOP20, DRVF14, AWEB22, AVC09, BMPB23, BDMP12, CSG^{+}24, CSLZ10, DDH23, DHC10, EGSM21, FS09, GR08, GCC06, GW24, HLLL21, LLL12, MZ07, MIRG06, MOP21, PBW^{+}20, PDCE11, SLJS08, SÇP24, WS10, XLX12, XGB^{+}25, FB15]. Batching [CGJZ15]. Bayesian [Mam11]. Behavior [DBDF19, FPC23, KTK^{+}16, PDCE11, RTH19, LDD06]. Behavior-based [PDCE11]. Behaviors [BBC^{+}11, DW15]. Behavioural [BEE^{+}20]. Benefit [WHA^{+}25]. Benefit-Cost-Risk-Aware [WHA^{+}25]. Best [BE16, CPS17, Edi14, FMVC14, PH15, SI17, VDK16]. Bike [CPZZ20]. Bike-sharing [CPZZ20]. Bilateral [CW14]. Billing [RB17]. Binary [GMMB15]. Bio [GMM12, FMS08].

Bio-Inspired [GMM12, FMS08]. biology [BCD^{+}06]. Bionic [DXP14]. Bird [MR11]. Birth [KD16]. Birth-Death [KD16]. BitTorrent [LXX^{+}14]. BizOps [FBL18]. block [GYSD08]. Blockchain [LTS^{+}25]. Blockchain-Based [LTS^{+}25]. bootstrapping [MT09]. boundary [GJM12]. bounding [SLJS08]. BPMN [SC19]. Budget [RB17, ZLHC21]. Budget-Driven [RB17]. Build [LNS24, RH16]. built [ZS09].

CAAC [VMG14]. Cache [LCT^{+}18]. Caching [XGB^{+}25]. Calibration [TKFL25]. Calibration-Free [TKFL25]. call [PDCE11]. Camera [LEC^{+}15, MM17]. Cameras [DAHS21]. Can [BEE^{+}20]. Capacity [WUK^{+}18]. Capital [PPB17]. Case [Buc19, GMM12]. Causal [FHDD23, PRJ11]. Cellular [MG11]. Centralized [ZCS12]. Centric [GGPTRC16, LVP15, MCPB20]. Certification [ST13]. Challenges [GCA^{+}24, PSA12, ST09]. Changing [KKK^{+}16, FFJ^{+}12]. characterization [GL08]. Characterizing [RTN^{+}17]. Checking [BBDB15, CSG^{+}24, SLW25, HSL^{+}07]. checkpoints [CY07]. Chemical [VCMZ11]. Chemical-Inspired [VCMZ11]. Choosing [LLL12]. Chowkidar [BLK^{+}09]. churn [LMSM12]. churn-resilient [LMSM12]. Circle [DLIP08]. class [KGJ08]. Classification [JH13, KIW06, PCL^{+}25]. Clock [TKFL25]. Cloud [AWEB22, FBL18, GB14, LCQB19, SMK21, TVKB16, WVT^{+}17, WUK^{+}18, ZLS22, ZSLG16]. Cloud-based [AWEB22]. Clouds [GS18, LHAES23, RB17]. Cluster [LHAES23]. Clustering [BMPB23, dASH16, GR08, QPGS12, ZCS12]. Clusters [LWQL16, dASH16, SA06]. Cluttered [KLWS16]. Co [YZS^{+}25]. Co-Optimization [YZS^{+}25]. Coalition [PBARA14, SZY^{+}20]. coevolution [WNV12b]. Cognitive [CMP13, MPC^{+}15]. Collaborative [FPC24, LSZ25, RH16, WLC^{+}25, LV07]. Collection [DZJ^{+}21]. Collective [Buc19, CHC16, HSC^{+}18, KKK^{+}18b, PPCE22, PPA18, SS12]. collectives [FSW^{+}10]. Collisionless [SBMM17]. Combining [CPZZ20, ZLW^{+}25]. Common [PBM14]. Common-Pool [PBM14]. Communication [XZL11, BDMP12, FSW^{+}10]. communications [DDF^{+}06]. Community [HWA^{+}20]. Compact [MLsRA^{+}15]. Comparison [MHP^{+}12]. compasses [SDY09]. Compensation [TKFL25]. Complex [BCC^{+}17, HEC^{+}16, ONC17, JI07]. Complexity [BEK09, CEK14]. Component [EYCM16]. Composed [SFdL25]. Composition [AOK11, Bak11, DBA^{+}21, HS11, MBB11, SHRB13, WCW^{+}17, FS09]. Comprehensive [BHQT24]. Computational [Gab11, VA11]. Computer [Mam11, DK12]. Computing [Bak11, BMS11, FRA^{+}25, GCA^{+}24, MHP^{+}12, MKS21, PSPR15, TWS20, BCD^{+}06, BCC^{+}12, HSM^{+}12, KGJ08, Lit07, SMSÇ^{+}10, TMC^{+}11, WBSI10]. Concepts [PSA12]. Conceptual [CGPP12]. Concurrency [LF19]. Conditional [ST13]. Conference [TWS20]. conflicts [DNT09]. Congestion [PLQ^{+}25]. Connected [CHZ^{+}25, LTS^{+}25, PLQ^{+}25, QKSV25, WLC^{+}25, XGB^{+}25]. Connection [RMKM17]. Connectivity [KRM16]. Consensus [BR11, GMMB15]. conserving [SLJS08]. Considerations [GS18]. Consistency [RTN^{+}17]. consistent [SDY09]. Consortium [LTS^{+}25]. Constant [JB11]. constrained [BMPB23, SLJS08]. Constraint [ZLHC21, MHZ13]. Constraints [SLW25, CY07].

Construction [BWO17, War19].

Consumption [FCD^{+}18]. Container [TOLG^{+}24]. Containerized [SMK21]. Containment [CLW^{+}14]. Content [XGB^{+}25, SA06]. Context [BCC^{+}23, FS09, WHH^{+}10b, WHH10a]. context-aware [FS09]. context-driven [WHH^{+}10b, WHH10a]. Contingent [MKS21]. Continuous [CW14, FP17, GEB^{+}21]. Continuous-Time [CW14]. Control [ALKG22, APSM18, ARS17, BGW24, BDMP12, CZW^{+}25, DAHS21, EK21, FSF24, FMA^{+}17, FDMD15, HWH^{+}17, KLWS16, KKK^{+}18a, KKK^{+}18b, LZ13, LF19, LDL16, MVV14, RMKM17, SWM19, SKA20, VMG14, WUK^{+}18, XLXZ14, GYSD08, KM08, LR12, LND12, WCD^{+}09]. Control-based [BDMP12]. Control-theoretic [SWM19]. Controllers [SMK21, SCC17, SFdL25]. Controlling [BWO17, EK21, KTK^{+}16, KKK^{+}16, KKK^{+}18a, RMM19]. Convention [HLLL21, VSMS13]. Convergence [FE12, KB12, PB13, ZSA09]. Cooperation [ACW10, PBARA14, TGT^{+}06]. Cooperative [ASS^{+}15, HLM15, MVV14, TMC^{+}11]. Coordination [FMVC14, GZBC25, HLM15, VCMZ11, MPBMP^{+}10]. correction [CLHX12]. correctness [HSL^{+}07]. Cost [BEE^{+}20, BWW^{+}17, GS18, KKK^{+}16, WHA^{+}25]. Coupled [LCT^{+}18, KB12]. Coverage [LDL16, PPCE22, GJM12]. Crash [BR11]. Creating [MSA09]. Credible [GZBC25]. cross [CSLZ10, PSB^{+}12]. cross-entropy [CSLZ10]. cross-layer [PSB^{+}12]. Crossing [HWH^{+}17]. Crowdsourcing [JAJ^{+}18, MPC^{+}15]. Cyber [CMS23, LVP15, SJN18]. Cyber-Physical [LVP15, CMS23].

Data [CMP13, DRPQ14, LVP15, LF19, LCT^{+}18, Men16, MCPB20, PRRR15, RDKB15, dASH16, ZSA09]. Data-Analytics [LF19]. Data-Centric [LVP15]. data-driven [ZSA09]. Database [GS18]. databases [SA06]. Deadline [ZCVL13]. Deadline-Driven [ZCVL13]. Dealing [GW24]. Death [KD16]. Decentralised [ONC17, PPCE22]. Decentralized [AOK11, ARS17, KGJ12, KB15, LND12, PPA18, RDKB15, SKA20, QPGS12]. Deception [Sar24]. Decision [AA16, GSB24, KKK^{+}18b, MHP^{+}12, MCGS18, WHA^{+}25, SS12]. Decision-Making [KKK^{+}18b, MHP^{+}12, MCGS18, WHA^{+}25]. Deep [BYH^{+}25, CHZ^{+}25, TSLF23, WGQV22]. deeper [XLX12]. Defending [LXX^{+}14]. Defense [DXP14, KD07]. Defined [HWH^{+}17, LNS24]. Defining [FP17]. Degree [JB11]. Delay [LZ13, SLJS08]. Delegation [FPC24]. Demand [BW19, GEB^{+}21]. Demonstration [BWO17]. dependable [DCL^{+}12]. Dependencies [EYCM16]. dependency [PRJ11]. dependent [MHZ13]. Deployed [ZSLG16]. Deployment [SZB19, WDTS11]. Description [Dua11]. Design [ARS17, BCD^{+}06, BBDB15, CHC16, CW11, CMP13, FE12, GDA10, GCA^{+}24, DW15, QPGS12, PPSM07]. Designing [LR12, WV18, War19, YHT16, ZSLG16]. Despite [BR11]. Detecting [DGL^{+}11]. Detection [CLW^{+}14, FRA^{+}25, IJDZ16, ONC17, TOLG^{+}24, SQX^{+}07, YTW08, ZS09]. detectors [ZS09]. Develop [BSK20]. Developing [LZC24]. Developments [Bah24]. Device [BVPD17, DY08]. Devices [BSK20, Das12]. DevOps [FBL18]. DIFC [SKA20]. different [APA12]. differential [APA12, CEA08]. dimensional [WCD^{+}09]. Disaster [SZB19]. Discovery [Bak11, CW11, Dua11, FGB11, DHC10]. Discrete [SMHP15]. Dispersion [Bea15]. disruption [XWN09]. disruption-tolerant [XWN09]. Dissemination [CMP13, MCPB20]. dissolution [VSMS13]. Distributed [BMPB23, BMS11, DGL^{+}11, FB15, FSW^{+}10, GMMB15, HMF^{+}15, KLWS16, LVP15, LEC^{+}15, MM17, MVV14, Men16, PRRR15, RPG^{+}15, RTN^{+}17, SHRB13, SMHP15, WVT^{+}17, War19, ZLW^{+}25, BCD^{+}06, Dat08, Dat09, HSL^{+}07, LMSM12, LR12, RYC^{+}07, SLJS08, WMA12, ZCS12]. Distribution [BVPD17, GB14]. Distributive [PBM14]. Disturbances [GMMB15]. Diverse [AH21, LDL16]. Diversification [AWEB22]. Diversity [LC21]. Diversity-driven [LC21]. Division [LDD06]. Domains [CW14]. downloading [DHJ08]. Drift [GW24]. Driven [BSS^{+}14, BBDB15, GGV20, RB17, VG14, ZCVL13, BW09, LC21, LHAES23, MIRG06, PSB^{+}12, WHH^{+}10b, WHH10a, ZSA09]. drivers [DY08]. Driving [BYH^{+}25]. Drone [PLQ^{+}25]. Dumb [KRM16]. DVFS [CGJZ15]. Dynamic [AWEB22, BEE^{+}20, CPZZ20, DBDF19, DBA^{+}21, GGV20, HBMA21, LEC^{+}15, MBB11, NCKA25, PBARA14, SC19, TSLF23, CY07, DCL^{+}12, FS09, SA06, USC^{+}08, WHH^{+}10b, WHH10a]. Dynamically [BGW24, FFJ^{+}12]. Dynamics [CZW^{+}25, XLXZ14, JI07, WNV12a].

e-Sampling [BWW^{+}17]. Economic [FBL18, PSA12]. Economies [PPA18]. Ecosystems [CMRZ15]. Edge [BHQT24, FRA^{+}25, GCA^{+}24, LHAES23, LTS^{+}25, XGB^{+}25, YZS^{+}25]. Editorial [Bah24, LV10, Nus18, PZ11, PZ13, PZ18, VP09, WBSI10]. Effective [VA11, WUK^{+}18, WGQV22]. Efficiency [CGJZ15, Das12]. Efficient [BYH^{+}25, CFGM16, DAHS21, FRA^{+}25, GYP12, HSC^{+}18, HLLL21, MCGS18, NCKA25, WGQV22, WXZ10]. Eigenspace [SQX^{+}07]. Elastic [DRPQ14, Men16, WXG^{+}25]. Elasticity [GS18]. Electronic [PPB17]. Embedded [BSK20, JH13, RYC^{+}07]. Emergence [HSC^{+}18, HLLL21, MALS22, ONC17, VSMS13]. Emergencies [VMG14]. Emergent [FP17, KTK^{+}16]. Empowered [FSW^{+}10]. Enabled [BNL24, YZS^{+}25, LC21, LCQB19]. enabling [CDV09]. encoding [MS12]. Energy [CGJZ15, LWQL16, SLJS08]. Enforcement [FSF24, RMM19]. Enforcing [CMS23]. Engineering [APSM18, CMRZ15, DDH23, ESBT19, VG14, PSFC12]. Enhancing [PRB16]. Enki [LC21]. Ensembles [Buc19]. enterprise [MSA09]. Entities [AOK11]. entropy [CSLZ10]. Entry [MAFS^{+}18]. Environment [CHCZ^{+}24, Gab11, PBW^{+}20]. Environments [BEE^{+}20, BCC^{+}17, DZJ^{+}21, FRA^{+}25, KLWS16, KKK^{+}16, MDC17, SHRB13, SZB19, VA11, ZSLG16, DCL^{+}12, DHC10, FFJ^{+}12, GPTW13, GDA10, Her10, LV07, MIRG06, TMC^{+}11]. Epidemic [XLXZ14, XLX12]. Equilibrium [CMS23, CEA08]. erasure [MS12]. erasure-resilient [MS12]. Establishment [SZB19]. Estimation [GEB^{+}21, ZSA09]. EUREMA [VG14]. Evaluation [CMP13, HBMA21, MCPB20, PAS^{+}25, DC12, GDA10, QPGS12]. Event [BWW^{+}17, HEC^{+}16, JH13, SMHP15, PRJ11]. Event-Sensitive [BWW^{+}17]. eventually [SDY09]. Evidence [MOP21, WS10]. Evidence-based [WS10]. E}volution [BCF^{+}08, SCC17, APA12, TMC^{+}11]. evolutionary [WDTS11]. evolvable [LS09]. Evolved [HWH^{+}17]. Evolving [BEE^{+}20, MHZ13]. exchange [Das12]. Exchanging [LCT^{+}18]. Execution [PRB16]. Executions [SJN18]. exercise [CDGT08]. Experience [BBC^{+}11].

Experiments [PSA12]. Explain [WHY24]. Explainable [MLFP24, WHY24]. Explanation [ALKG22]. Explicit [WUK^{+}18]. Exploiting [AHM09, HBDD14]. Expression [KIW06]. Extended [CPS17, PH15, SFdL25, VDK16]. Extending [PPSM07]. Extracting [VHK^{+}17]. Extraction [SC19].

Facial [KIW06]. Factorization [FG15]. factors [WNET07]. fair [Das12]. Farewell [PZ18]. Fast [CLW^{+}14, DP16, JH13, KKK^{+}16]. Fast-Spreading [CLW^{+}14]. Faster [SÇP24]. Fault [AD09, FG15, RYC^{+}07, WCD^{+}09]. fault-tolerant [WCD^{+}09]. faults [CLHX12]. Feature [BWO17]. featuring [FMS08]. Federated [ZLW^{+}25]. Feedback [RW20]. Feedforward [TKFL25]. Fewer [SÇP24]. Field [SPB24]. File [LCT^{+}18]. Filtering [LSZ25]. Filters [KCH14]. Finding [SZY^{+}20]. Fine [RB17]. Fine-Grained [RB17]. firewall [CLHX12]. First [CLHX12, Nus18]. Fishing [DXP14]. Flexible [AH21, MCGS18, SFdL25, CGPP12, MS12]. Flight [MR11]. Flight-Inspired [MR11]. Floating [PSPR15]. Floating-Point [PSPR15]. Flow [MAFS^{+}18, SKA20]. Flows [IJDZ16]. Foraging [DZJ^{+}21, LDD06]. Foreword [PK23, SCM24]. Formal [ARS17, CD11, DLPT14, DW15, BCC^{+}12, WMA12]. formalized [PSB^{+}12]. Formally [ZLS22]. Formation [BRII21, KLWS16, PBARA14, DLIP08, GJM12]. formations [GLMN09]. FORMS [WMA12]. forums [POPM07]. Fostering [PBARA14]. fragments [PSFC12]. Framework [AH21, ADAR25, BHQT24, BYH^{+}25, BDLM11, CSG^{+}24, CPZZ20, DDH23, FLHP24, FGB11, GEB^{+}21, LZC24, LZL^{+}24b, MS15, PTW07, PCL^{+}25, TOLG^{+}24, ZLHC21, AVC09, GJM12, LS09, WXZ10].

Free [TKFL25, SA12]. Function [LSZ25]. Functional [GSB24]. Functions [BHQT24]. fundamental [CDV09]. Fuzzy [LZ13, AGLV10].

Gabriel [MG11]. Game [LZL^{+}24b, Men16, RDKB15, YHT16, AVC09]. Game-Theoretic [Men16, RDKB15]. Game-Theoretical [LZL^{+}24b]. Games [CMGS16, AL09, CEA08]. Gap [HWH^{+}17]. gather [SDY09]. Gathering [SBMM17]. geared [WS10]. general [GL08]. generation [GR10]. Generative [LZL^{+}24a]. Generic [FDMD15, DNT09]. Genetic [DDH23, LNS24]. Geo [GS18]. Geo-Elasticity [GS18]. Geometric [BMS11]. Gesture [HMF^{+}15]. Gist [HLLL21]. GLDAP [TSLF23]. Global [TSLF23]. Goal [SZY^{+}20]. Goals [CV19, SC19, SDOP20]. Gossiping [DP16]. Governing [Sar24]. Grained [RB17]. Granularity [HBMA21]. Graph [HEC^{+}16, KTK^{+}16, RPG^{+}15, DKMD11]. Graphs [DBDF19, MG11, MKS21]. GraphStep [DKMD11]. Grid [BNL24, CY07, FMS08]. Grids [DRPQ14, Dua11, MG11, GYSD08]. G}roup [BCF^{+}08, ADV16, LDD06]. Groups [AMS^{+}19]. Growth [HWH^{+}17]. Guarantee [LZ13]. Guarantees [SWM19]. guided [CHZ^{+}25].

HAMLET [EGSM21]. Handle [SWM19]. hardware [DKMD11]. hash [LMSM12]. hash-tables [LMSM12]. Hazardous [DZJ^{+}21]. Healing [GGV20, MS15]. health [BLK^{+}09]. Help [BEE^{+}20]. Heterogeneity [LEC^{+}15, WNV12a]. Heterogeneous [EK21, FGB11, FDMD15, SHRB13, GDA10]. Heterogenous [YZS^{+}25]. Heuristic [HSC^{+}18, WDTS11]. heuristic/evolutionary [WDTS11]. Heuristics [CMP13]. Hierarchical [EGSM21, KKK^{+}18a, HSL^{+}07]. High [Dua11, PLQ^{+}25, PPSM07]. High-Availability [PLQ^{+}25]. High-Performance [Dua11]. Hoc [CW11, MR11, SZB19, PRJ11, SLJS08]. Holonic [FDMD15, HKR08]. Home [BDLM11]. Host [CLW^{+}14, SS12]. Host-Based [CLW^{+}14]. Human [CHCZ^{+}24, FPC23, FPC24, MCPB20, RW20]. Human-AI [FPC24]. Human-centric [MCPB20]. Human-machine [CHCZ^{+}24]. Hybrid [EK21, FPC24, Sar24, WDTS11]. Hyper [SCC17]. Hyper-Learning [SCC17]. hypernetwork [JI07]. Hypernetworks [ZLW^{+}25]. Hyperparameters [SÇP24].

IaaS [RB17]. ICT [BNL24]. ICT-Enabled [BNL24]. Identification [SÇP24]. IEC [FSF24]. IEEE [TS07]. III [POPM07]. Imitation [RH16]. immunologically [LS09]. immunologically-inspired [LS09]. Immunology [CHC16]. Impairment [RMKM17]. Implementation [CHC16, CW11, DKMD11, KM08]. implementations [BW09]. Improve [MVV14]. Improvement [CGJZ15, APA12]. Improving [APA12, FHDD23, GGV20, LF19, AHM09]. In-Memory [DRPQ14]. incentive [WNV12a]. Incentives [CPZZ20]. Including [AH21]. Increase [RDKB15]. incremental [GPTW13]. indulgence [GL08]. Industrial [ALKG22, FSF24]. Industry [WGA^{+}23]. Inference [BYH^{+}25]. Inferring [EYCM16]. Influence [RTH19]. Influence-aware [RTH19]. Information [KGL21, KKK^{+}18b, SKA20]. Informed [KB15]. infrared [KIW06]. infrastructure [SA12]. Infrastructureless [FMSA11]. Infrastructures [VMG14]. Inherently [MDC17]. inhibitory [KB12]. inhibitory-coupled [KB12]. Initial [KB15]. Innovative [SZB19]. insights [XLX12]. Inspired [GMM12, MR11, VCMZ11, XZL11, FMS08, GR10, KGJ08, LDD06, LS09]. Instances [PRB16]. Institutions [PSA12]. INSYTE [PCL^{+}25]. Integrals [KD16]. Integrated [YZS^{+}25]. Integrating [WCW^{+}17]. Intelligence [CLRC21, LZC24, WHY24, AGLV10, DHC10, Her10, LV10]. Intelligent [CW14, QKSV25, DHC10]. Intensity [VHK^{+}17]. Intensive [LZL^{+}24b]. Interaction [CZW^{+}25, EYCM16, MZ07, Pos07]. Interactions [HLLL21, RMM19]. Interactive [KM08, RW20]. Interdomain [VGR^{+}15]. Interface [BW19]. International [TWS20]. Internet [BVPD17, CGJZ15, USC^{+}08]. Interoperable [AGLV10, FGB11]. interpretation [KIW06]. Interventions [FHDD23]. Introduction [BCC^{+}12, BE16, BN12, Dat08, Dat09, DGB24, Edi14, LPZZ09, MCJ25, POPM07, QKSV25, SI17, Ser06, TWS20, TS07, ZP12]. Intrusion [IA18, SQX^{+}07, YTW08, ZS09]. invariant [HSL^{+}07]. ION [PBW^{+}20]. IoP [MCPB20]. IoT [AWEB22]. IoV [XGB^{+}25]. Isolation [MSA09]. Issue [Bak11, DGB24, MCJ25, PK23, QKSV25, SCM24, TWS20, Dat08, Dat09, LPZZ09, LV10, POPM07, TS07, VP09, WBSI10].

JADE [BBC^{+}11]. jointly [SLJS08]. Journal [Bah24]. Just [BRDA16]. Just-in-Time [BRDA16]. Justice [PBM14].

Kalman [KCH14]. Key [PRRR15, RTN^{+}17, WNV12a]. Key-Value [PRRR15]. keying [EGK08]. Knob [WUK^{+}18]. Knowledge [KPO19, FFJ^{+}12, MT09, MIRG06]. knowledge-driven [MIRG06].

labor [LDD06]. Laboratory [BCF^{+}08]. Landscape [ST09]. Language [DLPT14, SGP13]. Language-Level [SGP13]. Large [GGV20, KKK^{+}16, KKK^{+}18a, MCPB20, RPG^{+}15, WGQV22, AD09, WCD^{+}09]. Large-Scale [KKK^{+}16, KKK^{+}18a, RPG^{+}15, MCPB20, AD09, WCD^{+}09]. Largest [SZY^{+}20]. Latency [CMGS16, MCGS18, RTN^{+}17]. Latency-Aware [CMGS16, MCGS18]. layer [PSB^{+}12]. Lean [JH13]. Learning [BYH^{+}25, BSK20, CSG^{+}24, CHZ^{+}25, EGSM21, FHDD23, FP17, GPTW13, GF19, GWQ21, GW24, GZBC25, HL13, HLM15, HSC^{+}18, HLLL21, KB15, LC21, LCQB19, MDC17, MLFP24, MAFS^{+}18, PPA18, RW20, RTH19, SCC17, SÇP24, TSLF23, TOLG^{+}24, VGR^{+}15, WCW^{+}17, WGQV22, WLC^{+}25, XGB^{+}25, ZLHC21, ZLW^{+}25, XWN09]. Learning-based [CSG^{+}24, GW24]. Learning-enabled [LC21]. less [SDY09]. Level [SGP13]. Libraries [RMM19]. Lifelong [GW24]. Light [CDV09]. Lightweight [FE12, KKK^{+}16]. like [CSLZ10]. limited [SDY09]. Link [VGR^{+}15, ZSA09]. Literature [ESBT19, GWQ21]. Load [BGW24, GB14, JZL15, VHK^{+}17, AHM09, GYP12]. Local [LCT^{+}18]. Localization [MOP21, GCC06]. localizations [RYC^{+}07]. Locally [DGL^{+}11]. location [AHM09]. Logarithmic [EGK08]. Logistic [DBDF19]. Long [GZBC25]. Long-term [GZBC25]. Look [ZLS22]. Loosening [EK21]. Low [BWW^{+}17]. Low-Cost [BWW^{+}17]. Lyapunov [CHZ^{+}25]. Lyapunov-guided [CHZ^{+}25]. LyDRL [CHZ^{+}25].

M [ZS09]. M-AID [ZS09]. Machine [CSG^{+}24, EGSM21, GWQ21, TOLG^{+}24, XWN09, CHCZ^{+}24]. MACODO [WHH^{+}10b, WHH10a]. Macro [Mam11, BMZ12]. Maintenance [SDOP20, War19]. Making [AA16, BW19, GSB24, KKK^{+}18b, MHP^{+}12, MCGS18, WHA^{+}25]. Malware [DXP14]. managed [PPA18]. Management [ADAR25, BDLM11, BW19, HEC^{+}16, KPO19, LHAES23, MM17, MAFS^{+}18, PBM14, ZLS22]. Management-Based [MM17]. Managing [BHQT24, KGL21, LWQL16]. MANET [BDS07]. MANETs [XWN09]. Many [CMS25, MG11]. Many-objective [CMS25]. Map [ZHSP20]. MAPE [CHCZ^{+}24, DW15]. MAPE-K [CHCZ^{+}24, DW15]. MARC [FCD^{+}18]. Market [LCQB19]. Markets [TVKB16]. MAS [DRVF14]. Matching [LCQB19]. mathematical [WS10]. Matrix [FG15]. Matter [BMPB23]. Measure [AMS^{+}19]. measurement [KIW06]. measures [AD09]. Mechanism [KKK^{+}18b, TVKB16, DRVF14, CSLZ10]. Mechanisms [PAS^{+}25, RDKB15, APA12, BDS07, WNV12a]. Media [SR16]. Memory [DRPQ14, AL09, SDY09]. memory-less [SDY09]. message [BW09]. message-driven [BW09]. Metaverse [YZS^{+}25, ADAR25, QKSV25]. Metaverse-Integrated [YZS^{+}25]. Method [BBDB15, FE12, KKK^{+}18a, LDC^{+}18, LTS^{+}25, ZHSP20, CSLZ10, Lit07, PSFC12]. methodologies [PSFC12]. Methods [XZL11, BCC^{+}12, FSW^{+}10]. Metric [MG11]. Micro [SFdL25]. Micro-controllers [SFdL25]. Microgrids [FDMD15]. Microservice [HBMA21, WXG^{+}25, YZS^{+}25]. Microservices [BGW24]. Middleware [FGB11, FRL09, WHH^{+}10b, ZS09]. Migrations [LCT^{+}18, WVT^{+}17]. MIL [SÇP24]. MIL-based [SÇP24]. MiniMax [CEA08]. minority [AL09]. Misconfigurations [PAS^{+}25]. Mission [ZNU21]. Mitigating [BSS^{+}14]. Mix [LWQL16]. Mixing [Bea15]. Mobile [BCC^{+}17, CW11, MR11, SZB19, WVT^{+}17, XZL11, DLIP08, FRL09, SDY09]. Mobility [Buc19]. Mode [JB11]. Model [APSM18, BSS^{+}14, BBDB15, CMS25, CSG^{+}24, IA18, LCQB19, LHAES23, NCKA25, PBW^{+}20, SLW25, VG14, GCC06, WS10, WHH10a, WMA12]. Model-Based [IA18, PBW^{+}20]. Model-Driven [BSS^{+}14, VG14, LHAES23]. Modeling [ALKG22, BBDB15, DBDF19, DBA^{+}21, FCD^{+}18, FPC23, KD16, LXX^{+}14, MCPB20, SQX^{+}07, VHK^{+}17, WNV12a, ZCVL13, CGPP12, WNV12b]. Modelling [BCC^{+}23]. Models [BN12, FRA^{+}25, VA11]. Modular [BMPB23, LZC24]. Modulation [WUK^{+}18]. Monitoring [BWW^{+}17, CV19, FG15, ZHSP20, BLK^{+}09, HSL^{+}07]. morphogenetic [GJM12]. Motion [HWH^{+}17]. Multi [ADV16, AH21, Buc19, CLSS^{+}13, DDH23, GGPTRC16, GB14, GZBC25, LV07, MDC17, MALS22, PPB17, PPCE22, SJN18, WCW^{+}17, XGB^{+}25, ZLHC21, DHC10, DC12, HAMR13, HKR08, MHZ13, Pos07, TGT^{+}06, USC^{+}08, ZCS12]. Multi-Agent [ADV16, CLSS^{+}13, LV07, MDC17, PPB17, WCW^{+}17, XGB^{+}25, DDH23, GZBC25, MALS22, ZLHC21, HKR08, Pos07]. Multi-Agents [Buc19]. Multi-Cloud [GB14]. multi-constraint [MHZ13]. multi-objective [HAMR13]. multi-policy [DC12]. Multi-Robot [AH21, SJN18, PPCE22, TGT^{+}06, ZCS12]. multi-society-based [DHC10]. Multi-Tasking [AH21]. Multi-Tenant [GGPTRC16]. multi-tier [USC^{+}08]. Multiagent [HL13, HLM15, JAJ^{+}18, SQX^{+}07, WS10]. multicast [AVC09, SLJS08, XVYH11]. Multidimensional [GMM12]. Multilateral [HLLL21]. Multilayered [LV07]. multilevel [JI07]. Multimedia [MM17]. Multiobjective [FDMD15]. Multiplex [JZL15]. multirate [XVYH11]. Multirobot [DZJ^{+}21, KLWS16, GJM12]. Multiscale [FDMD15]. Multitolerance [CEK14]. mutation [WXZ10]. mute [BW09]. Mutual [RTH19].

Natural [HWH^{+}17]. nature [GR10, KGJ08]. nature-inspired [GR10, KGJ08]. necessary [CY07]. Negative [KTK^{+}16]. Negotiation [CW14, GZBC25, SR16, GR08, PTW07]. NEPTUNE [BHQT24]. Nervous [DXP14]. Nets [SLW25]. Network [Dua11, FE12, IJDZ16, SQX^{+}07, SZB19, BLK^{+}09, GSD08, LS09, LR12]. Networked [BWW^{+}17, CEA08]. networking [LPZZ09]. Networks [AMG18, BRII21, CW11, CMP13, FGB11, GMMB15, JZL15, KRM16, KKK^{+}16, KKK^{+}18a, LEC^{+}15, LNS24, LDL16, LXX^{+}14, MM17, Mam11, MR11, MOP21, MPC^{+}15, RMKM17, SSVB23, XLXZ14, YZS^{+}25, ACW10, AD09, DK12, FRL09, GLMN09, HSL^{+}07, LLL12, MPBMP^{+}10, MT09, MS12, PRJ11, SA12, VSMS13, WCD^{+}09, WNV12a, WNV12b, WNET07, XVYH11, XLX12, ZSA09]. Neural [LZ13]. Nodes [KRM16]. Non [GSB24, MDC17]. Non-Functional [GSB24]. Non-Stationary [MDC17]. Nonfunctional [FYC^{+}24]. Norm [MALS22]. Normative [MLsRA^{+}15]. Norms [ADV16, HSC^{+}18, SDOP20]. Number [dASH16].

objective [CMS25, HAMR13]. Observable [GSB24]. Obstacles [CV19]. Offloading [CHZ^{+}25, LTS^{+}25]. omega [BW09]. Ongoing [Bah24]. Online [IJDZ16, MKS21, MLFP24, MLsRA^{+}15, SCC17, TOLG^{+}24, QPGS12]. Open [ASS^{+}15, ST13, RYC^{+}07]. Operational [BNL24]. operators [WXZ10]. Opponents [CW14]. Opportunistic [BUL^{+}18, CMP13, MPC^{+}15]. Optimal [BW09, BR11, BRDA16, HL13, KKK^{+}18a, LND12]. optimistic [Das12]. Optimization [CPZZ20, LHKK23, LDC^{+}18, MHP^{+}12, YZS^{+}25, ZCVL13, DC12, HAMR13, WDTS11]. optimizer [WXZ10]. Optimizing [FPC24, SMK21, GYSD08, LR12]. Options [WV18]. Orchestration [SMHP15]. Ordering [SJN18]. organic [SMSÇ^{+}10, WBSI10]. Organisations [ADV16]. Organised [KPO19, PBM14]. Organising [PPB17]. Organization [AA16, PSPR15, DRVF14, CSLZ10, SMSÇ^{+}10, WHH10a]. organizations [KGJ12, WHH^{+}10b, WHH10a]. Organized [KKK^{+}16, GJM12, Her10, SSVB23]. Organizing [AOK11, KRM16, KKK^{+}18a, KKK^{+}18b, PSA12, TWS20, BMZ12, BDS07, FSW^{+}10, FMS08, KB12, LS09, LPZZ09, PRJ11, PSFC12, WCD^{+}09]. Oriented [DRVF14]. Oscar [GDA10]. oscillators [KB12]. Our [BMS11]. Outcomes [HL13]. Overlay [GMM12, GDA10, WNV12b, WNET07]. Overlays [JB11]. Overload [BGW24]. Overview [DC12].

P2P

[BDS07, CSLZ10, GMM12, JB11, LLL12]. P2P-like [CSLZ10]. Papers [BE16, CPS17, Edi14, PH15, SI17, TWS20, VDK16]. Parallel [MVV14]. Parallelization [CFGM16]. Parameter [BRDA16]. Parsimonious [GR10]. Partially [GSB24]. Particle [WXZ10]. partitioned [GYP12]. Partitioning [RPG^{+}15]. Partner [PBARA14]. partners [LLL12]. pattern [GJM12]. patterns [BCD^{+}06]. Peer [LXX^{+}14, DHJ08, HSM^{+}12, KGJ08, LMSM12, WNET07]. Peer-to-Peer [LXX^{+}14, DHJ08, HSM^{+}12, KGJ08, LMSM12, WNET07]. Percentile [LZ13]. Percentile-Based [LZ13]. Perception [FP17]. Performance [BSS^{+}14, CGJZ15, CMP13, Dua11, FHDD23, FRA^{+}25, GS18, LF19, LDC^{+}18, SMK21, SFdL25, ZCVL13, Lit07, MSA09]. performance-robust [MSA09]. Periods [RB17]. Persistence [TSLF23]. persistent [BDMP12]. Personalized [ZLW^{+}25].

Perspective [HWA^{+}20, JAJ^{+}18]. perturbations [GYP12]. Pervasive [Bak11, BDLM11, CMRZ15, CD11, Dua11, Gab11, MZ07, SHRB13, VCMZ11, BCC^{+}12, DC12, GPTW13, HSM^{+}12, SF12, ZP12]. Pheromone [ZHSP20, MZ07]. pheromone-based [MZ07]. philosophers [DNT09]. Physical [LVP15, SJN18, CMS23]. PID [SMK21]. Piecewise [LSZ25]. Pig [ZCVL13]. Pirates [ZHSP20]. Placement [BCC^{+}17, PRRR15, Her10]. Plan [MKS21]. Planning [FLHP24, IA18, KB15, LHKK23, MKS21, ZLS22]. Plans [SDOP20]. Plants [HWH^{+}17]. Plasticity [BEE^{+}20]. Platform [EGSM21]. playing [WNV12a]. Point [PSPR15]. points [MSA09]. Policies [BSK20, KB15]. Policy [GF19, SR16, CLHX12, DC12]. Pool [PBM14]. pools [LND12]. population [AAFJ08]. Power [BNL24, LR12]. Predicates [DGL^{+}11]. Predicting [FYC^{+}24, FPC23]. Prediction [CZW^{+}25, CLRC21, MDC17, ZLW^{+}25]. Prediction-Based [MDC17]. Predictive [APSM18, MVV14, XVYH11]. predictor [PDCE11]. Preferences [MBB11, SZY^{+}20, GPTW13]. Prescriptive [BBDB15]. Presence [SMHP15]. Preserving [BYH^{+}25]. Price [BR11]. Prices [VGR^{+}15]. Primate [XZL11]. Primate-Inspired [XZL11]. Principles [GCA^{+}24, PSA12]. Privacy [BYH^{+}25, SR16]. Privacy-Preserving [BYH^{+}25]. Proactive [MCGS18, VMG14]. Probabilistic [CV19, CSG^{+}24, GF19]. problem [GCC06]. problem-solving [GCC06]. Process [LCT^{+}18]. Processes [KD16, BW09]. Processing [HEC^{+}16, Men16, PSPR15]. Profiles [VHK^{+}17]. Profiling [LDC^{+}18]. Profitability [WUK^{+}18]. Programmable [BMPB23]. Programming [DLPT14, DDH23, HBDD14, LNS24, Mam11]. Programming-based [DDH23]. Programs [ZCVL13, BEK09]. Properties [BDS07]. Property [BBDB15]. Property-Driven [BBDB15]. Prosocial [MALS22]. Protecting [YEM14]. Protection [IA18]. Protocol [MR11, BDMP12]. Protocols [FE12, AAFJ08, CDV09, Pos07]. Providing [GS18]. Provisioning [GB14, KCH14, LZ13, SA06, USC^{+}08]. PSINES [CLRC21]. psychology [AVC09]. PTZ [DAHS21]. Public [HBDD14, WUK^{+}18]. pull [XLX12]. pull-based [XLX12]. Push [XLX12]. Push- [XLX12].

QoS [AHM09, DBA^{+}21, GSD08, ZLW^{+}25]. QoS-aware [DBA^{+}21]. QSTMF [ADAR25]. Quantum [ADAR25]. Quantum-Secured [ADAR25]. queries [GYP12].

Radio [BW19]. random [GYP12]. Randomization [SSVB23]. Range [MOP21]. Range-based [MOP21]. Ranking [WNET07]. Rational [VA11, ZS09]. Re [YZS^{+}25]. Re-Scheduling [YZS^{+}25]. REACT [PBW^{+}20]. REACT-ION [PBW^{+}20]. Reactive [SA06, WV18, GCC06]. reading [MS12]. Reality [HWH^{+}17]. Reasoning [SDOP20]. Rebalancing [CPZZ20]. Recognition [HMF^{+}15]. Recommendation [LSZ25]. reconfigurability [RYC^{+}07]. Reconfigurable [BSK20, PRJ11]. Reconfiguration [MVV14]. Reduction [WGQV22]. Reestablishment [KRM16]. reference [WMA12]. Reflective [SF12]. Regulation [CLSS^{+}13]. Reinforcement [BSK20, CHZ^{+}25, GF19, GZBC25, HL13, HLM15, KB15, MDC17, MLFP24, MAFS^{+}18, RW20, SÇP24, TSLF23, VGR^{+}15, WCW^{+}17, XGB^{+}25, ZLHC21]. reinforcing [VSMS13]. RelaxDHT [LMSM12]. Reliability [PRB16]. Reliable [BLK^{+}09, JZL15, AVC09]. Renewable [LWQL16]. Repair [War19]. Replicated [DRPQ14]. Replicating [FPC23]. Replication [PRB16, LMSM12]. Requirement [FYC^{+}24]. Requirements [APSM18, GSB24]. Research [GCA^{+}24, HWA^{+}20, LZL^{+}24a, ST09]. Residual [ZLW^{+}25]. Resilience [CMS23]. Resilient [BRII21, LMSM12, MS12]. Resolution [CV19]. Resource [ASS^{+}15, FCD^{+}18, GEB^{+}21, JH13, KCH14, KPO19, LF19, LHAES23, PBM14, SSN^{+}12, SMHP15, ZLS22, LND12]. Resource-Lean [JH13]. Resources [YZS^{+}25, AL09]. Response [IA18, ZS09]. results [BEK09, PB13]. retrieval [MIRG06]. Reuse [GF19, KGL21]. Review [ESBT19, GWQ21, GCA^{+}24]. Reviewers [Ano06, Ano07, Ano08, ACM06, Ano09]. Revised [CPS17, PH15, VDK16]. revising [BEK09]. Reward [GGV20]. Rewriting [HEC^{+}16]. RFID [MZ07, XZL11]. Rigorously [DW15]. Risk [DZJ^{+}21, WHA^{+}25]. Risk-aware [DZJ^{+}21]. Roadmap [LZL^{+}24a, WHY24]. Robot [AH21, BMPB23, BBDB15, KD16, SJN18, SBMM17, SCC17, WV18, War19, GLMN09, JI07, PPCE22, TGT^{+}06, ZCS12]. Robot-based [BMPB23]. robots [DLIP08, LDD06, SDY09]. Robust [CLSS^{+}13, HSC^{+}18, LC21, VSMS13, MSA09]. robustness [KB12]. role [RYC^{+}07, WNV12a]. roles [RYC^{+}07]. Routing [MR11, VGR^{+}15, MHZ13, PRJ11, PB13, SLJS08, ZSA09]. rtual [BCF^{+}08]. rule [GR10]. run [HSL^{+}07]. run-time [HSL^{+}07]. Runtime [CV19, CMS23, FSF24, KTK^{+}16, NCKA25, PBW^{+}20, RMM19, RTH19].

SAC [FMVC14]. Safe [DHJ08, GF19]. Safety [ST13, Dat08, Dat09]. Sampling [BWW^{+}17, LVP15]. SAPERE [CMRZ15]. SARDE [GEB^{+}21]. SASO [CPS17, PH15, VDK16]. Satisfaction [GSB24]. Scalability [DBA^{+}21, FHDD23, GGV20]. Scalable [FBL18, JB11, PRRR15, ZLS22, BLK^{+}09]. Scale [KKK^{+}16, KKK^{+}18a, RPG^{+}15, AD09, MCPB20, WCD^{+}09]. Scaler [DRPQ14]. Scaling [DRPQ14, WXG^{+}25]. ScatterD [WDTS11]. SCEL [DLPT14]. Scenarios [AH21]. Scheduling [PLQ^{+}25, RB17, YZS^{+}25]. Scientific [RB17, HAMR13]. SDN [MAFS^{+}18]. SEAMS [BE16, Edi14, MCJ25, SI17, SCM24]. Search [KGL21]. secret [SA12, MOP21]. Section [BE16, Edi14, SI17, BCC^{+}12, BN12, ZP12]. Secure [ALKG22, BYH^{+}25, MOP21, WLC^{+}25]. Secured [ADAR25]. Securing [AWEB22, LZL^{+}24b]. Security [TOLG^{+}24, Dat08, Dat09, SA12]. SeDiM [FGB11]. Selected [CPS17, PH15, VDK16, TWS20]. Selection [Gab11, HS11, SSN^{+}12, CY07, DHC10, SS12]. Self [AA16, AOK11, APSM18, AAFJ08, ARS17, BVPD17, BMZ12, BBC^{+}11, CMGS16, CMS25, CSG^{+}24, CGJZ15, DXP14, DNT09, DY08, DP16, ESBT19, FB15, FMVC14, FCD^{+}18, FP17, FMA^{+}17, GSB24, GGV20, GWQ21, GW24, GLMN09, GEB^{+}21, Her10, HWA^{+}20, HEC^{+}16, DW15, KRM16, KGL21, KB12, KPO19, KKK^{+}16, KKK^{+}18a, KKK^{+}18b, LZ13, LHKK23, LCQB19, LZL^{+}24a, LZL^{+}24b, LNS24, MS15, MHP^{+}12, MCGS18, PRRR15, PSPR15, PRJ11, PPB17, PSA12, PBM14, PPA18, PAS^{+}25, RMKM17, RTH19, ST09, SGP13, Sar24, SSVB23, SWM19, SPB24, TWS20, TOLG^{+}24, DRVF14, VG14, WCD^{+}09, WGQV22, WGA^{+}23, WHA^{+}25, WLC^{+}25, YHT16, YEM14, ACW10, BDS07, BN12, CSLZ10, DHJ08, FSW^{+}10, FRL09, FMS08, GYSD08, GR10, GJM12, KGJ08, KGJ12, LS09, LPZZ09, PSFC12, SMSÇ^{+}10, TGT^{+}06, VSMS13, WMA12]. Self-* [KGL21]. Self-Adaptation [BVPD17, CMGS16, FMVC14, GSB24, GW24, LZL^{+}24b, MCGS18, RMKM17, WGA^{+}23, CMS25, RTH19, KGJ12]. Self-adapting [CSG^{+}24]. Self-Adaptive [AA16, APSM18, ARS17, FMA^{+}17, GW24, GEB^{+}21, DW15, LZ13, LZL^{+}24a, PAS^{+}25, SGP13, SPB24, VG14, WHA^{+}25, WLC^{+}25, GWQ21, HWA^{+}20, ST09, SWM19, WGQV22, BN12, FRL09, KGJ08, LPZZ09, WMA12]. Self-Adaptiveness [PSPR15]. Self-Adaptivity [LNS24]. Self-Assembly [FP17, GR10, TGT^{+}06]. Self-Aware [FCD^{+}18, LHKK23]. Self-awareness [ESBT19]. Self-Defense [DXP14]. self-downloading [DHJ08]. Self-Governing [Sar24]. Self-Healing [GGV20, MS15]. Self-Learning [LCQB19]. Self-managed [PPA18]. Self-Management [HEC^{+}16]. Self-Optimization [MHP^{+}12]. self-optimizing [GYSD08]. Self-Organised [KPO19, PBM14]. Self-Organising [PPB17]. Self-Organization [PSPR15, DRVF14, CSLZ10, SMSÇ^{+}10]. Self-Organized [KKK^{+}16, Her10, SSVB23, GJM12]. Self-Organizing [AOK11, KRM16, KKK^{+}18a, KKK^{+}18b, PSA12, TWS20, BMZ12, KB12, PRJ11, WCD^{+}09, BDS07, FSW^{+}10, FMS08, LS09, LPZZ09, PSFC12]. Self-Protecting [YEM14]. self-reconfigurable [PRJ11]. self-reinforcing [VSMS13]. Self-Self [BBC^{+}11]. self-similar [ACW10]. Self-Stabilized [DP16]. Self-Stabilizing [FB15, YHT16, AAFJ08, DNT09, DY08, GLMN09]. Self-Supervised [TOLG^{+}24]. Self-Tuning [CGJZ15, PRRR15]. selfish [CDGT08]. semantic [GR08]. Semantics [FS09]. Semantics-based [FS09]. Semi [DDH23]. Semi-automated [DDH23]. Sensing [BWW^{+}17]. Sensitive [BWW^{+}17].

Sensor [AMG18, BCC^{+}17, BUL^{+}18,

HMF^{+}15, KRM16, LDL16, MM17, MOP21, RMKM17, AD09, BLK^{+}09, FRL09, HSL^{+}07, MPBMP^{+}10, ZSA09]. Sensors [JH13, XZL11, BMZ12]. SEPPDL [BYH^{+}25]. Sequential [FG15]. Server [LWQL16, SA06]. Serverless [BHQT24]. Servers [CGJZ15, KCH14]. Service [AOK11, Bak11, CMRZ15, Dua11, DBA^{+}21, FCD^{+}18, FGB11, HS11, LCQB19, SSN^{+}12, DRVF14, WCW^{+}17, WVT^{+}17, FS09, GYSD08, Her10, MIRG06, PTW07]. Service-Oriented [DRVF14]. Services [BNL24, CW11, GGPTRC16, VCMZ11, AGLV10, TMC^{+}11]. Setting [BRDA16]. Shared [ZLW^{+}25, LND12, SA12, SA06]. shared-secret [SA12]. Sharing [BUL^{+}18, PPA18, ZLHC21, CPZZ20]. SHõ [MS15]. SimCA* [SWM19]. similar [ACW10]. S}imulation [BCF^{+}08]. Simulations [CMGS16, SMHP15]. Simultaneous [FLHP24]. Situated [LV07]. Situation [PBW^{+}20]. Situation-aware [PBW^{+}20]. Size [BMPB23]. Size-constrained [BMPB23]. skin [KIW06]. Small [CHCZ^{+}24, JB11, PAS^{+}25]. Smart [BW19, Buc19, FDMD15, LEC^{+}15, VMG14, GSD08]. Smartphone [BW19]. SMT [FB15]. SMT-Based [FB15]. snap [CDV09]. snap-stabilization [CDV09]. So-Grid [FMS08]. soccer [JI07]. Social [BCF^{+}08, HL13, HLM15, PPB17, SR16, ACW10, AVC09, VSMS13]. Socially [HL13]. Societies [Sar24]. society [DHC10]. Socio [KD07, PSA12]. Socio-Economic [PSA12]. Socio-technical [KD07]. SOD [BW19]. Software [AMG18, APSM18, ESBT19, EYCM16, FP17, FMA^{+}17, LZL^{+}24b, LNS24, SMK21, SGP13, VG14, YEM14, MIRG06, PPSM07, ST09]. Software-Defined [LNS24]. Software-Intensive [LZL^{+}24b]. solving [GCC06]. Space [MG11]. Spaces [GW24, VCMZ11, WGQV22]. spamming [KD07]. sparse [DKMD11]. Spatial [BMS11, DKMD11, FMSA11, Mam11, VCMZ11, WDTS11]. Spatiotemporal [HMF^{+}15]. Special [Bak11, BE16, DGB24, Edi14, LV10, MCJ25, PK23, QKSV25, SI17, SCM24, TWS20, TS07, WBSI10, BCC^{+}12, BN12, Dat08, Dat09, LPZZ09, POPM07, VP09, ZP12]. Specification [CD11, WMA12]. Specifying [Pos07]. Spot [PRB16, TVKB16]. Spreading [CLW^{+}14, XLX12]. Spyware [DXP14]. Stabilisability [BRII21]. Stability [MVV14, ZSA09]. stabilization [CDGT08, CDV09, Dat08, Dat09]. Stabilized [DP16]. Stabilizing [FB15, YHT16, AAFJ08, DNT09, DY08, GLMN09]. Stable [CHZ^{+}25]. State [Bah24, LZL^{+}24a, MHZ13]. state-dependent [MHZ13]. States [BNL24]. Static [LEC^{+}15, XZL11]. station [LR12]. Stationary [MDC17]. Status [WXG^{+}25]. Status-aware [WXG^{+}25]. StatuScale [WXG^{+}25]. Stealth [JB11]. Steering [HWH^{+}17]. Steiner [SLJS08]. step [CLHX12]. Stepwise [LDC^{+}18]. Stochastic [CMGS16, KGL21, PB13, ZCS12]. Storage [FMSA11, RDKB15, MS12]. Stores [PRRR15, RTN^{+}17]. Strategies [DZJ^{+}21, FMA^{+}17, IA18, LHKK23, MHP^{+}12, HAMR13, WNV12b]. Strategy [WXG^{+}25, LMSM12]. Stream [Men16]. Streaming [LDC^{+}18]. Streams [dASH16]. Strict [SZY^{+}20]. Structural [DRVF14]. structure [WNV12b]. Structured [GDA10]. structures [HAMR13, VSMS13]. Study [MLFP24, SFdL25]. Sub [SQX^{+}07]. Sub-Eigenspace [SQX^{+}07]. Subgoal [SÇP24]. Subnetworks [AWEB22]. Successful [SZY^{+}20]. sufficient [CY07]. Superdiffusive [Bea15]. Supervised [TOLG^{+}24]. Support [EYCM16, SGP13, dASH16, HSM^{+}12]. Supporting [DCL^{+}12, SC19, RYC^{+}07].

Survey

[FPC23, WGA^{+}23, YEM14, DDF^{+}06].

Sustainability [FBL18]. SUTC’06 [TS07]. Swarm [CFGM16, ZHSP20, WXZ10]. swarming [LR12]. Swarms [Bea15, BBDB15, BWO17, EK21, KD16, PLQ^{+}25, PSPR15, WV18]. Switching [PBARA14]. Synchronisation [TKFL25]. Synchronization [SSVB23, KB12]. synergizing [APA12]. Synthesis [FB15, MLsRA^{+}15]. System [AWEB22, CV19, DAHS21, Gab11, HSM^{+}12, LV07, SJN18, dASH16, MS12, YTW08]. Systematic [DBA^{+}21, ESBT19, GWQ21, YEM14]. Systems [ALKG22, ASS^{+}15, APSM18, ARS17, Bah24, BMS11, BWW^{+}17, BCC^{+}23, CMS23, CLSS^{+}13, CHC16, CSG^{+}24, CHCZ^{+}24, CMP13, CD11, DXP14, DLPT14, DBDF19, DDH23, FB15, FYC^{+}24, FSF24, FMA^{+}17, GWQ21, GW24, GMM12, HL13, HLM15, HWA^{+}20, DW15, JAJ^{+}18, KGL21, LC21, LVP15, LZL^{+}24a, LZL^{+}24b, MHP^{+}12, MALS22, MLFP24, MLsRA^{+}15, ONC17, PPB17, PPCE22, PCL^{+}25, PAS^{+}25, RDKB15, SMK21, ST13, SWM19, TWS20, WGQV22, WHA^{+}25, WLC^{+}25, YHT16, YEM14, BDS07, BN12, CY07, Dat08, Dat09, DC12, HKR08, JI07, KM08, LPZZ09, Lit07, MSA09, Pos07, PSFC12, RYC^{+}07, SMSÇ^{+}10, SF12, SQX^{+}07, TGT^{+}06, WS10, WMA12].

TAAS [Bah24, Ano09]. Tables

[SLW25, LMSM12]. tabu [WXZ10]. Tags [XZL11, MZ07]. Take [BMS11]. Task [AH21, BRII21, CHZ^{+}25, FLHP24, JZL15, LTS^{+}25, MBB11, PRB16, SJN18, ZCS12]. Tasking [AH21]. taxonomy [PSB^{+}12]. taxonomy-driven [PSB^{+}12]. Team [BRII21]. Teaming [CHCZ^{+}24]. Teams [FPC24, War19, ZCS12]. technical [KD07, POPM07]. Techniques [FFJ^{+}12, WCW^{+}17]. Temperature [TKFL25, KIW06]. Templates [DW15]. temporal [CY07, CGPP12, GPTW13]. Temporary [RMKM17]. Tenant [GGPTRC16]. term [GZBC25]. Test [LC21]. testbeds [BLK^{+}09]. Testing [SPB24]. their [MG11]. Theoretic [Men16, RDKB15, SWM19]. Theoretical [LZL^{+}24b, AVC09, GYSD08]. Theory [MOP21, YHT16, KM08]. Things [BVPD17]. Three [GB14, WCD^{+}09]. three-dimensional [WCD^{+}09]. Three-Tier [GB14]. Thresholds [XLX12, XLXZ14]. Tier [GB14, USC^{+}08]. Tight [SBMM17]. Time [BRDA16, CW14, HSL^{+}07, MHZ13]. tolerance [AD09]. tolerant [WCD^{+}09, XWN09]. Topology [LDL16, MM17, RMKM17, MT09, WCD^{+}09]. Trace [HLLL21]. Trace-based [HLLL21]. Tracking [KLWS16, GCC06]. Tradeoff [RTN^{+}17]. Traditional [PCL^{+}25]. Traffic [CZW^{+}25, FSW^{+}10]. Train [LC21]. Training [FRA^{+}25]. Trajectory [CZW^{+}25]. Transactional [DRPQ14, DRPQ14]. Transactions [Bah24, DK12]. transfer [GYSD08]. Transparent [CFGM16]. Transportation [HBDD14]. tree [SLJS08]. Tropos [PPSM07]. Trust [AA16, ADAR25, BNL24, LCQB19, SKA20, VA11, WS10]. Trust-Aware [SKA20]. Trust-Based [AA16]. Trust-enabled [LCQB19]. Trusted [LZC24, Das12]. Trustworthy [HS11]. TSLAM [LCQB19]. Tuning [CGJZ15, PRRR15, SÇP24, YTW08]. Tuple [VCMZ11].

UAVs [ZHSP20, ZNU21]. Ubiquitous [Bak11, CD11, Dua11, LV07, TMC^{+}11]. Uncertainty [CMS25, FLHP24, HWA^{+}20, KGL21, KKK^{+}18b, SWM19, SMHP15]. Uncrewed [PAS^{+}25]. Understanding [BCC^{+}23, JAJ^{+}18]. Underwater [LDL16, MOP21]. unified [WXZ10]. Unifying [WMA12]. Units [LF19].

UNITY [BEK09]. Unknown

[CLW^{+}14, CW14]. Unmanned [CHCZ^{+}24]. unreliable [GLMN09]. Update [Bah24]. upon [ZS09]. Urban [Buc19, HBDD14]. Urban-Area [HBDD14]. Usage [VA11]. use [AL09]. User [CPZZ20, GGPTRC16, HWH^{+}17, MLFP24, MBB11, AHM09, GPTW13]. User-Centric [GGPTRC16]. User-Defined [HWH^{+}17]. users [GSD08]. Using [BSS^{+}14, BSK20, CMGS16, FP17, GW24, KCH14, KD16, LNS24, MAFS^{+}18, PRB16, RH16, SMK21, SSVB23, SDY09, YHT16, ZHSP20, Das12, FHDD23, HAMR13, HSL^{+}07, KIW06, MKS21]. Utility [GGV20, DRVF14]. Utility-Based [DRVF14]. Utility-Driven [GGV20].

Value [AMS^{+}19, PRRR15, RTN^{+}17]. values [ZLHC21]. VANETs [ADAR25]. variability [PPSM07]. Variable [dASH16]. variations [KIW06]. Vehicle [CZW^{+}25, MPBMP^{+}10]. Vehicles [CHZ^{+}25, LTS^{+}25, PLQ^{+}25, QKSV25, WLC^{+}25, XGB^{+}25]. Verification [ARS17, CMS23, CD11, NCKA25, CY07]. Verified [ZLS22]. Very [JB11]. Vi} [BCF^{+}08]. Via [LF19, CMS23, CSG^{+}24, FSF24, PB13, ZCS12]. Viable [WV18]. Viewpoint [WHA^{+}25]. Violations [FYC^{+}24]. virtual [BMZ12]. Virtualized [KCH14]. Virus [DXP14]. ViSAGE [BCF^{+}08]. visibility [SDY09]. Vision [Bah24]. Visual [BWO17]. voice [KD07]. Volatility [BEE^{+}20]. vs [ZHSP20].

WA [MS15]. weak [DLIP08]. Web [ADAR25, GYSD08, MS15, PTW07, QKSV25, YZS^{+}25]. Web-Based [MS15]. Weighting [LSZ25]. Welcome [Bah24]. Whom [WHY24]. Wireless [AMG18, LDL16, MM17, RMKM17, SSVB23, SZB19, TKFL25, AHM09, AD09, BLK^{+}09, FSW^{+}10, FRL09, HSL^{+}07, LPZZ09, MPBMP^{+}10, MIRG06, SA12, WCD^{+}09]. within [SJN18]. WLANs [AHM09]. Workflow [PRB16, SLW25, CY07, HAMR13]. Workflows [RB17, SC19, CGPP12]. Workloads [LHAES23]. World [BMS11]. Worm [CLW^{+}14]. Worms [LXX^{+}14]. writing [MS12]. WSNs [BRDA16].

XtreemOS [SSN^{+}12].

References

Ahmadi:2016:TBD

[AA16] Kamilia Ahmadi and Vicki H.

Allan. Trust-based decision making in a self-adaptive agent organization. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11 (2):10:1–10:??, July 2016. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Angluin:2008:SSP

[AAFJ08] Dana Angluin, James Aspnes, Michael J. Fischer, and Hong Jiang. Self-stabilizing population protocols. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(4):13:1–13:??, November 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

TAAS-Staff:2006:R

[ACM06] ACM Transactions on Autonomous and Adaptive Systems staff. Reviewers. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1(1):114, September 2006. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Allen:2010:CTS

[ACW10] Stuart M. Allen, Gualtiero Colombo, and Roger M. Whitaker. Cooperation through self-similar social networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(1):4:1–4:??, February 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Ammari:2009:FTM

[AD09] Habib M. Ammari and Sajal K. Das. Fault tolerance measures for largescale wireless sensor networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(1):2:1–2:??, January 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Awan:2025:QQS

[ADAR25] Kamran Ahmad Awan, Ikram Ud Din, Ahmad Almogren, and Joel J. P. C. Rodrigues. QSTMF: Quantum-secured trust management framework for VANETs in Web 3.0 and Metaverse. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(3):21:1– 21:29, September 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Aldewereld:2016:GNM

[ADV16] Huib Aldewereld, Virginia Dignum, and Wamberto W. Vasconcelos. Group norms for multi-agent organisations. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(2):15:1– 15:??, July 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Acampora:2010:IAF

[AGLV10] Giovanni Acampora, Matteo Gaeta, Vincenzo Loia, and Athanasios V. Vasilakos. Interoperable and adaptive fuzzy services for ambient intelligence applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(2):8:1–8:??, May 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Arif:2021:FFD

[AH21] Muhammad Usman Arif and Sajjad Haider. A flexible framework for diverse multirobot task allocation scenarios including multi-tasking. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 16(1):3:1–3:??, March 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502200.

Alyfantis:2009:EUL

[AHM09] George Alyfantis, Stathes Hadjiefthymiades, and Lazaros Merakos. Exploiting user location for load balancing WLANs and improving wireless QoS. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(2):13:1–13:??, May 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Araujo:2009:UMR

[AL09] Ricardo M. Araujo and Luis C. Lamb. On the use of memory and resources in minority games. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(2):11:1– 11:??, May 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Adepu:2022:MAE

[ALKG22] Sridhar Adepu, Nianyu Li, Eunsuk Kang, and David Garlan. Modeling and analysis of explanation for secure industrial control systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 17(3–4):5:1– 5:??, December 2022. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3557898.

Afanasov:2018:SAW

[AMG18] Mikhail Afanasov, Luca Mottola, and Carlo Ghezzi. Software adaptation in wireless sensor networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12 (4):18:1–18:??, January 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Alshebli:2019:MAV

[AMS^{+}19] Bedoor K. Alshebli, Tomasz P.

Michalak, Oskar Skibski, Michael Wooldridge, and Talal Rahwan. A measure of added value in groups. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(4):18:1– 18:??, July 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3335547.

Anonymous:2006:R

[Ano06]  Anonymous.  Reviewers.

ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1(2):260–261, December 2006. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Anonymous:2007:R

[Ano07] Anonymous. Reviewers 2007.

ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(4):17:1–17:??, November 2007. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Anonymous:2008:R

[Ano08] Anonymous. Reviewers 2008.

ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(4):21:1–21:??, November 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Anonymous:2009:TR

[Ano09] Anonymous. TAAS reviewers 2009. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(4):25:1– 25:??, November 2009. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Al-Oqily:2011:DSO

[AOK11] Ibrahim Al-Oqily and Ahmed Karmouch. A decentralized self-organizing service composition for autonomic entities. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(1):7:1–7:??, February 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Ali:2012:IDE

[APA12] Musrrat Ali, Millie Pant, and Ajith Abraham. Improving differential evolution algorithm by synergizing different improvement mechanisms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(2):20:1–20:??, July 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Angelopoulos:2018:ESA

[APSM18] Konstantinos Angelopoulos, Alessandro V. Papadopoulos, Vítor E. Silva Souza, and John Mylopoulos. Engineering self-adaptive software systems: From requirements to model predictive control. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(1):1:1–1:??, May 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Arcaini:2017:FDV

[ARS17] Paolo Arcaini, Elvinia Riccobene, and Patrizia Scandurra. Formal design and verification of self-adaptive systems with decentralized control. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(4):25:1– 25:??, February 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Anders:2015:CRA

[ASS^{+}15] Gerrit Anders, Alexander Schiendorfer, Florian Siefert, Jan-Philipp Steghöfer, and Wolfgang Reif. Cooperative resource allocation in open systems of systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10 (2):11:1–11:??, June 2015. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Anastasopoulos:2009:AFR

[AVC09] Markos P. Anastasopoulos, Athanasios V. Vasilakos, and Panayotis G. Cottis. An autonomic framework for reliable multicast: a game theoretical approach based on social psychology. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(4):21:1– 21:??, November 2009. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Almohri:2022:DSD

[AWEB22] Hussain Almohri, Layne Watson, David Evans, and Stephen Billups. Dynamic system diversification for securing cloud-based IoT subnetworks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 17(1–2):2:1– 2:??, June 2022. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547350.

Bahsoon:2024:ATA

[Bah24] Rami Bahsoon. ACM Transactions on Autonomous and Adaptive Systems (ACM TAAS): Editorial welcome and update on state of the journal, vision and ongoing developments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(2):8:1– 8:??, June 2024. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3661314.

Bakhouya:2011:SIA

[Bak11] Mohamed Bakhouya. Special issue: Adaptive service discovery and composition in ubiquitous and pervasive computing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(1):1:1–1:??, February 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bouchenak:2011:ASS

[BBC^{+}11] Sara Bouchenak, Fabienne Boyer, Benoit Claudel, Noel De Palma, Olivier Gruber, and Sylvain Sicard. From autonomic to self-self behaviors: The JADE experience. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(4):28:1–28:??, October 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Brambilla:2015:PDD

[BBDB15] Manuele Brambilla, Arne Brutschy, Marco Dorigo, and Mauro Birattari. Propertydriven design for robot swarms: a design method based on prescriptive modeling and model checking. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(4):17:1– 17:??, January 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bakhouya:2012:ISS

[BCC^{+}12] Mohamed Bakhouya, Roy Campbell, Antonio Coronato, Giuseppe de Pietro, and Anand Ranganathan. Introduction to special section on formal methods in pervasive computing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1): 6:1–6:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bartolini:2017:AMS

[BCC^{+}17] Novella Bartolini, Tiziana Calamoneri, Stefano Ciavarella, Thomas La Porta, and Simone Silvestri. Autonomous mobile sensor placement in complex environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12 (2):7:1–7:??, May 2017. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bumiller:2023:UCM

[BCC^{+}23] Anne Bumiller, Stéphanie Challita, Benoit Combemale, Olivier Barais, Nicolas Aillery, and Gael Le Lan. On understanding context modelling for adaptive authentication systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(1): 3:1–3:??, March 2023. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3582696.

Babaoglu:2006:DPB

[BCD^{+}06] Ozalp Babaoglu, Geoffrey Canright, Andreas Deutsch, Gianni A. Di Caro, Frederick Ducatelle, Luca M. Gambardella, Niloy Ganguly, Márk Jelasity, Roberto Montemanni, Alberto Montresor, and Tore Urnes. Design patterns from biology for distributed computing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1(1):26–66, September 2006. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Baumes:2008:VVR

[BCF^{+}08] Jeffrey Baumes, Hung-Ching (Justin)

Chen, Matthew Francisco, Mark Goldberg, Malik Magdon-Ismail, and William Wallace. ViSAGE: a vi rtual laboratory for simulation and analysis of social group evolution. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(3):8:1–8:??, August 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bourcier:2011:AAM

[BDLM11] Johann Bourcier, Ada Diaconescu, Philippe Lalanda, and Julie A. McCann. AutoHome: An autonomic management framework for pervasive home applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(1):8:1–8:??, February 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Blanchini:2012:CBP

[BDMP12] Franco Blanchini, Daniele De Caneva, Pier Luca Montessoro, and Davide Pierattoni. Control-based ppersistent adaptive communication protocol. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7 (2):29:1–29:??, July 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Biskupski:2007:PMS

[BDS07] Bartosz Biskupski, Jim Dowling, and Jan Sacha. Properties and mechanisms of selforganizing MANET and P2P systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(1): 1:1–1:??, March 2007. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bencomo:2016:ISS

[BE16] Nelly Bencomo and Gregor Engels. Introduction to the special section on best papers from SEAMS 2014. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(4):22:1– 22:??, February 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Beal:2015:SDM

[Bea15]  Jacob Beal.  Superdiffu-

sive dispersion and mixing of swarms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(2):10:1– 10:??, June 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Barnes:2020:BPC

[BEE^{+}20] Chloe M. Barnes, Anikó Ekárt, Kai Olav Ellefsen, Kyrre Glette, Peter R. Lewis, and Jim Tørresen. Behavioural plasticity can help evolving agents in dynamic environments but at the cost of volatility. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(4):11:1– 11:26, December 2020. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3487918.

Bonakdarpour:2009:CRR

[BEK09] Borzoo Bonakdarpour, Ali Ebnenasir, and Sandeep S. Kulkarni. Complexity results in revising UNITY programs. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(1):5:1–5:??, January 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bhattacharya:2024:DBL

[BGW24] Ratnadeep Bhattacharya, Yuan Gao, and Timothy Wood. Dynamically balancing load with overload control for microservices. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(4):22:1– 22:??, December 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3676167.

Baresi:2024:NCF

[BHQT24] Luciano Baresi, Davide Yi Xian Hu, Giovanni Quattrocchi, and Luca Terracciano. NEPTUNE: a comprehensive framework for managing serverless functions at the edge. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19 (1):7:1–7:??, March 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3634750.

Bapat:2009:CRS

[BLK^{+}09] S. Bapat, W. Leal, T. Kwon, P. Wei, and A. Arora. Chowkidar: Reliable and scalable health monitoring for wireless sensor network testbeds. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(1):3:1–3:??, January 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bassil:2023:DSC

[BMPB23] Jad Bassil, Abdallah Makhoul, Benoît Piranda, and Julien Bourgeois. Distributed sizeconstrained clustering algorithm for modular robotbased programmable matter. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(1):1:1–1:??, March 2023. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580282.

Beal:2011:SCD

[BMS11] Jacob Beal, Olivier Michel, and Ulrik Pagh Schultz. Spatial computing: Distributed systems that take advantage of our geometric world. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(2):11:1–11:??, June 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bicocchi:2012:SOV

[BMZ12] Nicola Bicocchi, Marco Mamei, and Franco Zambonelli. Selforganizing virtual macro sensors. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1): 2:1–2:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bouchachia:2012:ISS

[BN12] Abdelhamid Bouchachia and Nadia Nedjah. Introduction to the special section on selfadaptive systems: Models and algorithms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):13:1– 13:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Brand:2024:ATO

[BNL24] Michael Brand, Anand Narayan, and Sebastian Lehnhoff. Applying trust for operational states of ICT-enabled power grid services. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(4):25:1–25:??, December 2024. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3654672.

Bonnet:2011:PAO

[BR11] François Bonnet and Michel Raynal. The price of anonymity: Optimal consensus despite asynchrony, crash, and anonymity. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6 (4):23:1–23:??, October 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Brienza:2016:JTA

[BRDA16] Simone Brienza, Manuel Roveri, Domenico De Guglielmo, and Giuseppe Anastasi. Just-in-time adaptive algorithm for optimal parameter setting in 802.15.4 WSNs. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10 (4):27:1–27:??, February 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Barambones:2021:RTF

[BRII21] Jose Barambones, Florian Richoux, Ricardo Imbert, and Katsumi Inoue. Resilient team formation with stabilisability of agent networks for task allocation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(3):7:1– 7:24, September 2021. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3463368.

Burger:2020:RED

[BSK20] Alwyn Burger, Gregor Schiele, and David W. King. Reconfigurable embedded devices using reinforcement learning to develop action policies. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(4):14:1– 14:25, December 2020. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3487920.

Barna:2014:MAU

[BSS^{+}14] Cornel Barna, Mark Shtern, Michael Smit, Vassilios Tzerpos, and Marin Litoiu. Mitigating DoS attacks using performance model-driven adaptive algorithms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9 (1):3:1–3:??, March 2014. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bucchiarone:2019:CAT

[Buc19] Antonio Bucchiarone. Collective adaptation through multi-agents ensembles: The case of smart urban mobility. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(2):6:1–6:??, October 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Beal:2018:AOA

[BUL^{+}18] Jacob Beal, Kyle Usbeck, Joseph Loyall, Mason Rowe, and James Metzler. Adaptive opportunistic airborne sensor sharing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(1):6:1– 6:??, May 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Beal:2017:SAD

[BVPD17] Jacob Beal, Mirko Viroli, Danilo Pianini, and Ferruccio Damiani. Self-adaptation to device distribution in the Internet of Things. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12 (3):12:1–12:??, October 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Biely:2009:OMD

[BW09] Martin Biely and Josef Widder. Optimal messagedriven implementations of omega with mute processes. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(1):4:1–4:??, January 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Brocanelli:2019:SMS

[BW19] Marco Brocanelli and Xiaorui Wang. SOD: Making Smartphone smart on demand with radio interface management. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13 (3):12:1–12:??, March 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3275521.

Budhraja:2017:FCC

[BWO17] Karan K. Budhraja, John Winder, and Tim Oates. Feature construction for controlling swarms by visual demonstration. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(2):10:1– 10:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bhuiyan:2017:SES

[BWW^{+}17] Md Zakirul Alam Bhuiyan, Jie Wu, Guojun Wang, Tian Wang, and Mohammad Mehedi Hassan. esampling: Event-sensitive autonomous adaptive sensing and low-cost monitoring in networked sensing systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(1):1:1–1:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Bobo:2025:SSE

[BYH^{+}25] Wang Bobo, Hongwei Yang, Meng Hao, Jiannan Zhang, Hui He, and Weizhe Zhang. SEPPDL: a secure and efficient privacy-preserving deep learning inference framework for autonomous driving. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(1):5:1–5:??, March 2025. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3708505.

Coronato:2011:FSV

[CD11] Antonio Coronato and Giuseppe De Pietro. Formal specification and verification of ubiquitous and pervasive systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(1):9:1–9:??, February 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Cohen:2008:ESS

[CDGT08] Johanne Cohen, Anurag Dasgupta, Sukumar Ghosh, and Sébastien Tixeuil. An exercise in selfish stabilization. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(4):15:1–15:??, November 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Cournier:2009:LES

[CDV09] Alain Cournier, Stephane Devismes, and Vincent Villain. Light enabling snapstabilization of fundamental protocols. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(1): 6:1–6:??, January 2009. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Cao:2008:MEN

[CEA08] Hui Cao, Emre Ertin, and Anish Arora. MiniMax equilibrium of networked differential games. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(4):14:1– 14:??, November 2008. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Chen:2014:CAM

[CEK14] Jingshu Chen, Ali Ebnenasir, and Sandeep Kulkarni. The complexity of adding multitolerance. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(3):15:1– 15:??, October 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Cicirelli:2016:TEP

[CFGM16] Franco Cicirelli, Agostino Forestiero, Andrea Giordano, and Carlo Mastroianni. Transparent and efficient parallelization of swarm algorithms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(2):14:1– 14:??, July 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Cheng:2015:STB

[CGJZ15] Dazhao Cheng, Yanfei Guo, Changjun Jiang, and Xiaobo Zhou. Self-tuning batching with DVFS for performance improvement and energy efficiency in Internet servers. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(1):6:1–6:??, March 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Combi:2012:CMF

[CGPP12] Carlo Combi, Matteo Gozzi, Roberto Posenato, and Giuseppe Pozzi. Conceptual modeling of flexible temporal workflows. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(2):19:1–19:??, July 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Capodieci:2016:AIC

[CHC16] Nicola Capodieci, Emma Hart, and Giacomo Cabri. Artificial immunology for collective adaptive systems design and implementation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11 (2):6:1–6:??, July 2016. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Cleland-Huang:2024:HMT

[CHCZ^{+}24] Jane Cleland-Huang, Theodore Chambers, Sebastian Zudaire, Muhammed Tawfiq Chowdhury, Ankit Agrawal, and Michael Vierhauser. Humanmachine teaming with small unmanned aerial systems in a MAPE-K environment. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(1):3:1–3:??, March 2024. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3618001.

Chen:2025:LLG

[CHZ^{+}25] Yanming Chen, Ziyang Huang, Yiwen Zhang, Weiwei Fang, and Neal N. Xiong. Ly-DRL: Lyapunov-guided deep reinforcement learning for stable task offloading in connected autonomous vehicles. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(3):24:1– 24:29, September 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Chen:2012:FST

[CLHX12] Fei Chen, Alex X. Liu, Jeehyun Hwang, and Tao Xie. First step towards automatic correction of firewall policy faults. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(2):27:1–27:??, July 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Cumin:2021:PAA

[CLRC21] Julien Cumin, Grégoire Lefebvre, Fano Ramparany, and James L. Crowley. PSINES: Activity and availability prediction for adaptive ambient intelligence. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15 (1):1:1–1:12, February 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3424344.

Campos:2013:RRA

[CLSS^{+}13] Jordi Campos, Maite Lopez-Sanchez, Maria Salamó, Pedro Avila, and Juan A. Rodríguez-Aguilar. Robust regulation adaptation in multi-agent systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(3):13:1– 13:??, September 2013. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Chen:2014:HBA

[CLW^{+}14] Songqing Chen, Lei Liu, Xinyuan Wang, Xinwen Zhang, and Zhao Zhang. A hostbased approach for unknown fast-spreading worm detection and containment. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8 (4):21:1–21:??, January 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Camara:2016:ALA

[CMGS16] Javier Cámara, Gabriel A. Moreno, David Garlan, and Bradley Schmerl. Analyzing latency-aware self-adaptation using stochastic games and simulations. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10 (4):23:1–23:??, February 2016.

CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Conti:2013:DPE

[CMP13] Marco Conti, Matteo Mordacchini, and Andrea Passarella. Design and performance evaluation of data dissemination systems for opportunistic networks based on cognitive heuristics. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(3):12:1– 12:??, September 2013. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Castelli:2015:EPS

[CMRZ15] Gabriella Castelli, Marco Mamei, Alberto Rosi, and Franco Zambonelli. Engineering pervasive service ecosystems: The SAPERE approach. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(1):1:1–1:??, March 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Camilli:2023:ERC

[CMS23] Matteo Camilli, Raffaela Mirandola, and Patrizia Scandurra. Enforcing resilience in cyber-physical systems via equilibrium verification at runtime. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(3):12:1– 12:??, September 2023. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3584364.

Camilli:2025:MOS

[CMS25] Matteo Camilli, Raffaela Mirandola, and Patrizia Scandurra. Many-objective selfadaptation under model uncertainty. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(2):12:1– 12:28, June 2025. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Cabri:2017:SSR

[CPS17] Giacomo Cabri, Gauthier Picard, and Niranjan Suri. SASO 2016: Selected, revised, and extended best papers. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(3):11:1– 11:??, October 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Chiariotti:2020:BSO

[CPZZ20] Federico Chiariotti, Chiara Pielli, Andrea Zanella, and Michele Zorzi. A bikesharing optimization framework combining dynamic rebalancing and user incentives. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(3): 11:1–11:30, March 2020. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3376923.

Casimiro:2024:SAM

[CSG^{+}24] Maria Casimiro, Diogo Soares, David Garlan, Luís Rodrigues, and Paolo Romano. Selfadapting machine learningbased systems via a probabilistic model checking framework. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(3):18:1– 18:??, September 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3648682.

Chen:2010:SOM

[CSLZ10] Gang Chen, Abdolhossein Sarrafzadeh, Chor Ping Low, and Liang Zhang. A selforganization mechanism based on cross-entropy method for P2P-like applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(4):15:1–15:??, November 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Cailliau:2019:RMR

[CV19] Antoine Cailliau and Axel Van Lamsweerde. Runtime monitoring and resolution of probabilistic obstacles to system goals. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14 (1):3:1–3:??, September 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3337800.

Chen:2011:DIA

[CW11] Shyr-Kuen Chen and Pi-Chung Wang. Design and implementation of an anycast services discovery in mobile ad hoc networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6 (1):2:1–2:??, February 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Chen:2014:IAB

[CW14] Siqi Chen and Gerhard Weiss. An intelligent agent for bilateral negotiation with unknown opponents in continuous-time domains. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(3):16:1– 16:??, October 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Chen:2007:ASN

[CY07] Jinjun Chen and Yun Yang. Adaptive selection of necessary and sufficient checkpoints for dynamic verification of temporal constraints in grid workflow systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2 (2):6:1–6:??, June 2007. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Chen:2025:VDI

[CZW^{+}25] Jian Chen, Shaorui Zhou, Wei Wang, Yuzhu Hu, Jianqing Li, Ben guo He, Junxin Chen, Marwan Omar, Ali Kashif Bashir, and Xiping Hu. Vehicle dynamics and interaction for trajectory prediction and traffic control. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(2):11:1– 11:19, June 2025. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Davani:2021:ASE

[DAHS21] Sina G. Davani, Musab S. Al-Hadrusi, and Nabil J. Sarhan. An autonomous system for efficient control of PTZ cameras. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 16(2):6:1–6:??, June 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3507658.

Dashti:2012:EOF

[Das12] Mohammad Torabi Dashti. Efficiency of optimistic fair exchange using trusted devices. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):3:1–3:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Silva:2016:SSC

[dASH16] Jonathan de Andrade Silva and Eduardo Raul Hruschka. A support system for clustering data streams with a variable number of clusters.

ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(2):11:1– 11:??, July 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Datta:2008:ISI

[Dat08] Ajoy K. Datta. Introduction to special issue on stabilization, safety, and security of distributed systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(4):12:1–12:??, November 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Datta:2009:ISI

[Dat09] Ajoy K. Datta. Introduction to special issue on stabilization, safety, and security of distributed systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(1):1:1–1:??, January 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Duboc:2021:SSM

[DBA^{+}21] Leticia Duboc, Rami Bahsoon, Faisal Alrebeish, Carlos Mera-Gómez, Vivek Nallur, Rick Kazman, Philip Bianco, Ali Babar, and Rajkumar Buyya. Systematic scalability modeling of QoS-aware dynamic service composition. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 16(3–4):10:1–10:??, December 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3529162.

Demare:2019:ABM

[DBDF19] Thibaut Démare, Cyrille Bertelle, Antoine Dutot, and Dominique Fournier. Adaptive behavior modeling in logistic systems with agents and dynamic graphs. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13 (3):15:1–15:??, March 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3313799.

Dusparic:2012:AMP

[DC12] Ivana Dusparic and Vinny Cahill. Autonomic multipolicy optimization in pervasive systems: Overview and evaluation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):11:1– 11:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Dixit:2012:ASA

[DCL^{+}12] Mônica Dixit, António Casimiro, Paolo Lollini, Andrea Bondavalli, and Paulo Verissimo. Adaptare: Supporting automatic and dependable adaptation in dynamic environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(2):18:1–18:??, July 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Dobson:2006:SAC

[DDF^{+}06] Simon Dobson, Spyros Denazis, Antonio Fernández, Dominique Gaïti, Erol Gelenbe, Fabio Massacci, Paddy Nixon, Fabrice Saffre, Nikita Schmidt, and Franco Zambonelli. A survey of autonomic communications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1(2): 223–259, December 2006. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Donnell:2023:GPB

[DDH23] Nicola Mc Donnell, Jim Duggan, and Enda Howley. A genetic programming-based framework for semi-automated multi-agent systems engineering. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(2):6:1–6:??, June 2023. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3584731.

DiNitto:2024:IAS

[DGB24] Elisabetta Di Nitto, Ilias Gerostathopoulos, and Kirstie Bellman. Introduction to ACSOS 2022 special issue. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(3):14:1– 14:??, September 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3676168.

DeRosa:2011:DLD

[DGL^{+}11] Michael De Rosa, Seth Copen Goldstein, Peter Lee, Jason Campbell, and Padmanabhan S. Pillai. Detecting locally distributed predicates. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(2):13:1–13:??, June 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Duman:2010:MSB

[DHC10] Hakan Duman, Hani Hagras, and Victor Callaghan. A multi-society-based intelligent association discovery and selection for ambient intelligence environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5 (2):7:1–7:??, May 2010. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Dastidar:2008:SPP

[DHJ08] Kajari Ghosh Dastidar, Ted Herman, and Colette Johnen. Safe peer-to-peer self-downloading. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(4):19:1–19:??, November 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Dolev:2012:ATC

[DK12] Shlomi Dolev and Marina Kopeetsky. Anonymous transactions in computer networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(2):26:1–26:??, July 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Delorimier:2011:SHI

[DKMD11] Michael Delorimier, Nachiket Kapre, Nikil Mehta, and André Dehon. Spatial hardware implementation for sparse graph algorithms in Graph-Step. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(3):17:1– 17:??, September 2011. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Dieudonne:2008:CFW

[DLIP08] Yoann Dieudonné, Ouiddad Labbani-Igbida, and Franck Petit. Circle formation of weak mobile robots. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(4):16:1–16:??, November 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

DeNicola:2014:FAA

[DLPT14] Rocco De Nicola, Michele Loreti, Rosario Pugliese, and Francesco Tiezzi. A formal approach to autonomic systems programming: The SCEL language. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(2): 7:1–7:??, July 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Danturi:2009:SSP

[DNT09] Praveen Danturi, Mikhail Nesterenko, and Sébastien Tixeuil. Self-stabilizing philosophers with generic conflicts. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(1):7:1–7:??, January 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Dulman:2016:SSF

[DP16] Stefan Dulman and Eric Pauwels. Self-stabilized fast gossiping algorithms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(4):29:1–29:??, February 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Didona:2014:TAS

[DRPQ14] Diego Didona, Paolo Romano, Sebastiano Peluso, and Francesco Quaglia. Transactional Auto Scaler: Elastic scaling of replicated inmemory transactional data grids. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(2):11:1– 11:??, July 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Val:2014:UBM

[DRVF14] Elena Del Val, Miguel Rebollo, Mateo Vasirani, and Alberto Fernández. Utility-based mechanism for structural self-organization in serviceoriented MAS. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9 (3):12:1–12:??, October 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Duan:2011:NSD

[Dua11] Qiang Duan. Network service description and discovery for high-performance ubiquitous and pervasive grids. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(1):3:1–3:??, February 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Iglesia:2015:MKF

[DW15] Didac Gil De La Iglesia and Danny Weyns. MAPE-K formal templates to rigorously design behaviors for self-adaptive systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(3):15:1– 15:??, October 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Dai:2014:BAN

[DXP14] Y. S. Dai, Y. P. Xiang, and Y. Pan. Bionic autonomic nervous systems for self-defense against DoS, spyware, malware, virus, and fishing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(1):4:1–4:??, March 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Dolev:2008:SSD

[DY08] Shlomi Dolev and Reuven Yagel. Self-stabilizing device drivers. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(4):17:1– 17:??, November 2008. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Di:2021:RAC

[DZJ^{+}21] Kai Di, Yifeng Zhou, Jiuchuan Jiang, Fuhan Yan, Shaofu Yang, and Yichuan Jiang. Risk-aware collection strategies for multirobot foraging in hazardous environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 16(3–4):8:1– 8:??, December 2021. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3514251.

Editors:2014:ISS

[Edi14]  Editors.  Introduction to

the special section on best papers from SEAMS 2012. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(4):16:1–16:??, January 2014. CODEN ????

ISSN 1556-4665 (print), 1556-4703 (electronic).

Elmallah:2008:LK

[EGK08] Ehab S. Elmallah, Mohamed G. Gouda, and Sandeep S. Kulkarni. Logarithmic keying. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(4):18:1–18:??, November 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Esmaeili:2021:HHA

[EGSM21] Ahmad Esmaeili, John C. Gallagher, John A. Springer, and Eric T. Matson. HAM-LET: a hierarchical agentbased machine learning platform. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 16(3–4):9:1– 9:??, December 2021. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3530191.

Esterle:2021:LCH

[EK21] Lukas Esterle and David W. King. Loosening control — a hybrid approach to controlling heterogeneous swarms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 16(2):5:1–5:??, June 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3502725.

Elhabbash:2019:SAS

[ESBT19] Abdessalam Elhabbash, Maria Salama, Rami Bahsoon, and Peter Tino. Self-awareness in software engineering: a systematic literature review. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(2):5:1–5:??, October 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Esfahani:2016:ISC

[EYCM16] Naeem Esfahani, Eric Yuan, Kyle R. Canavera, and Sam Malek. Inferring software component interaction dependencies for adaptation support. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(4):26:1– 26:??, February 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Faghih:2015:SBS

[FB15] Fathiyeh Faghih and Borzoo Bonakdarpour. SMT-Based synthesis of distributed selfstabilizing systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(3):21:1–21:??, October 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Fokaefs:2018:DBE

[FBL18] Marios Fokaefs, Cornel Barna, and Marin Litoiu. From DevOps to BizOps: Economic sustainability for scalable cloud applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(4):25:1–25:??, January 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Ferroni:2018:MRC

[FCD^{+}18] Matteo Ferroni, Andrea Corna, Andrea Damiani, Rolando Brondolin, John D. Kubiatowicz, Donatella Sciuto, and Marco D. Santambrogio. MARC: a resource consumption modeling service for self-aware autonomous agents. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(4):21:1– 21:??, January 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Frey:2015:GHC

[FDMD15] Sylvain Frey, Ada Diaconescu, David Menga, and Isabelle Demeure. A generic holonic control architecture for heterogeneous multiscale and multiobjective smart microgrids. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(2):9:1–9:??, June 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Farahat:2012:LMA

[FE12] Aly Farahat and Ali Ebnenasir. A lightweight method for automated design of convergence in network protocols. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(4):38:1–38:??, December 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Fisch:2012:TKA

[FFJ^{+}12] Dominik Fisch, Dominik Fisch, Martin Jänicke, Edgar Kalkowski, and Bernhard Sick. Techniques for knowledge acquisition in dynamically changing environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):16:1– 16:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Feng:2015:FMS

[FG15] Dawei Feng and Cecile Germain. Fault monitoring with sequential matrix factorization. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(3):20:1– 20:??, October 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Flores:2011:SMF

[FGB11] Carlos Flores, Paul Grace, and Gordon S. Blair. SeDiM: a middleware framework for interoperable service discovery in heterogeneous networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(1):6:1–6:??, February 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Fadiga:2023:ICL

[FHDD23] Kanvaly Fadiga, Etienne Houzé, Ada Diaconescu, and Jean-Louis Dessalles. Improving causal learning scalability and performance using aggregates and interventions. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(3):11:1– 11:??, September 2023. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3607872.

Faruq:2024:FST

[FLHP24] Fatma Faruq, Bruno Lacerda, Nick Hawes, and David Parker. A framework for simultaneous task allocation and planning under uncertainty. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(4):21:1– 21:??, December 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3665499.

Filieri:2017:CSS

[FMA^{+}17] Antonio Filieri, Martina Maggio, Konstantinos Angelopoulos, Nicolás D’Ippolito, Ilias Gerostathopoulos, Andreas Berndt Hempel, Henry Hoffmann, Pooyan Jamshidi, Evangelia Kalyvianaki, Cristian Klein, Filip Krikava, Sasa Misailovic, Alessandro V. Papadopoulos, Suprio Ray, Amir M. Sharifloo, Stepan Shevtsov, Mateusz Ujma, and Thomas Vogel. Control strategies for self-adaptive software systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(4):24:1– 24:??, February 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Forestiero:2008:GSO

[FMS08] Agostino Forestiero, Carlo Mastroianni, and Giandomenico Spezzano. So-Grid: a self-organizing Grid featuring bio-inspired algorithms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(2):5:1–5:??, May 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Fernandez-Marquez:2011:ISS

[FMSA11] Jose Luis Fernandez-Marquez, Giovanna Di Marzo Serugendo, and Josep Lluis Arcos. Infrastructureless spatial storage algorithms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(2):15:1– 15:??, June 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Fernandez-Marquez:2014:BAS

[FMVC14] Jose Luis Fernandez-Marquez, Mirko Viroli, and Gabriella Castelli. Best ACM SAC articles on coordination and self-adaptation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9 (2):6:1–6:??, July 2014. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Filho:2017:DES

[FP17] Roberto Rodrigues Filho and Barry Porter. Defining emergent software using continuous self-assembly, perception, and learning. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(3):16:1– 16:??, October 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Fuchs:2023:MRP

[FPC23] Andrew Fuchs, Andrea Passarella, and Marco Conti. Modeling, replicating, and predicting human behavior: a survey. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(2):4:1–4:??, June 2023. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3580492.

Fuchs:2024:ODC

[FPC24] Andrew Fuchs, Andrea Passarella, and Marco Conti. Optimizing delegation in collaborative human-AI hybrid teams. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(4):23:1– 23:??, December 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3687130.

Fernando:2025:ETA

[FRA^{+}25] Duneesha Fernando, Maria A.

Rodriguez, Patricia Arroba, Leila Ismail, and Rajkumar Buyya. Efficient training approaches for performance anomaly detection models in edge computing environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(2):13:1– 13:27, June 2025. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Fok:2009:AMA

[FRL09] Chien-Liang Fok, Gruia-Catalin Roman, and Chenyang Lu. Agilla: a mobile agent middleware for self-adaptive wireless sensor networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4 (3):16:1–16:??, July 2009. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Fujii:2009:SBC

[FS09] Keita Fujii and Tatsuya Suda.

Semantics-based context-aware dynamic service composition. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(2):12:1–12:??, May 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Faqrizal:2024:AIC

[FSF24] Irman Faqrizal, Gwen Salaün, and Yliès Falcone. Adaptive industrial control systems via IEC 61499 and runtime enforcement. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(4):24:1– 24:??, December 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3691345.

Fekete:2010:EWC

[FSW^{+}10] Sándor P. Fekete, Christiane Schmidt, Axel Wegener, Horst Hellbrück, and Stefan Fischer. Empowered by wireless communication: Distributed methods for self-organizing traffic collectives. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(3):11:1– 11:??, September 2010. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Fang:2024:PNR

[FYC^{+}24] Xinwei Fang, Sinem Getir Yaman, Radu Calinescu, Julie Wilson, and Colin Paterson. Predicting nonfunctional requirement violations in autonomous systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(1):6:1–6:??, March 2024. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3632405.

Gaber:2011:ASA

[Gab11] Jaafar Gaber. Action selection algorithms for autonomous system in pervasive environment: a computational approach. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(1):10:1– 10:??, February 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Grozev:2014:MCP

[GB14] Nikolay Grozev and Rajkumar Buyya. Multi-cloud provisioning and load distribution for three-tier applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(3):13:1–13:??, October 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Golpayegani:2024:AEC

[GCA^{+}24] Fateneh Golpayegani, Nanxi Chen, Nima Afraz, Eric Gyamfi, Abdollah Malekjafarian, Dominik Schäfer, and Christian Krupitzer. Adaptation in edge computing: a review on design principles and research challenges. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(3):19:1– 19:??, September 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3664200.

Gechter:2006:RAB

[GCC06] Franck Gechter, Vincent Chevrier, and François Charpillet. A reactive agent-based problemsolving model: Application to localization and tracking. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1(2):189–222, December 2006. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Girdzijauskas:2010:SOH

[GDA10] Šarūnas Girdzijauskas, Anwitaman Datta, and Karl Aberer. Structured overlay for heterogeneous environments: Design and evaluation of Oscar. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(1):2:1–2:??, February 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Grohmann:2021:SFC

[GEB^{+}21] Johannes Grohmann, Simon Eismann, André Bauer, Simon Spinner, Johannes Blum, Nikolas Herbst, and Samuel Kounev. SARDE: a framework for continuous and selfadaptive resource demand estimation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(2):6:1– 6:31, June 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3463369.

Garcia:2019:PPR

[GF19] Javier García and Fernando Fernández. Probabilistic policy reuse for safe reinforcement learning. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13 (3):14:1–14:??, March 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3310090.

Garcia-Galan:2016:UCA

[GGPTRC16] Jesús García-Galán, Liliana Pasquale, Pablo Trinidad, and Antonio Ruiz-Cortés. User-centric adaptation analysis of multi-tenant services. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(4):24:1– 24:??, February 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Ghahremani:2020:ISR

[GGV20] Sona Ghahremani, Holger Giese, and Thomas Vogel. Improving scalability and reward of utility-driven self-healing for large dynamic architectures. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(3): 12:1–12:41, March 2020. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3380965.

Guo:2012:MFS

[GJM12] Hongliang Guo, Yaochu Jin, and Yan Meng. A morphogenetic framework for selforganized multirobot pattern formation and boundary coverage. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):15:1–15:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Guerraoui:2008:GCI

[GL08] R. Guerraoui and N. Lynch. A general characterization of indulgence. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(4):20:1– 20:??, November 2008. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Gilbert:2009:SSR

[GLMN09] Seth Gilbert, Nancy Lynch, Sayan Mitra, and Tina Nolte. Self-stabilizing robot formations over unreliable networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(3):17:1–17:??, July 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Giordanelli:2012:BIP

[GMM12] Raffaele Giordanelli, Carlo Mastroianni, and Michela Meo. Bio-inspired P2P systems: The case of multidimensional overlay. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(4):35:1–35:??, December 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Gogolev:2015:DBC

[GMMB15] Alexander Gogolev, Nikolaj Marchenko, Lucio Marcenaro, and Christian Bettstetter. Distributed binary consensus in networks with disturbances. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(3):19:1– 19:??, October 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Gallacher:2013:LUP

[GPTW13] Sarah Gallacher, Eliza Papadopoulou, Nick K. Taylor, and M. Howard Williams. Learning user preferences for adaptive pervasive environments: an incremental and temporal approach. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(1):5:1–5:??, April 2013. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Garruzzo:2008:ACB

[GR08] Salvatore Garruzzo and Domenico Rosaci. Agent clustering based on semantic negotiation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(2):7:1–7:??, May 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Grushin:2010:PRG

[GR10] Alexander Grushin and James A.

Reggia. Parsimonious rule generation for a nature-inspired approach to self-assembly. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(3):12:1–12:??, September 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Guo:2018:PCC

[GS18] Tian Guo and Prashant Shenoy. Performance and cost considerations for providing geo-elasticity in database clouds. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(4):19:1– 19:??, January 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Garcia:2024:DMS

[GSB24] Luis Garcia, Huma Samin, and Nelly Bencomo. Decision making for self-adaptation based on partially observable satisfaction of non-functional requirements. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19 (2):11:1–11:??, June 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3643889.

Gelenbe:2008:AQA

[GSD08] Erol Gelenbe, Georgia Sakellari, and Maurizio D’Arienzo. Admission of QoS aware users in a smart network. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(1):4:1–4:??, March 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Gheibi:2024:DDA

[GW24] Omid Gheibi and Danny Weyns. Dealing with drift of adaptation spaces in learningbased self-adaptive systems using lifelong self-adaptation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(1):5:1–5:??, March 2024. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3636428.

Gheibi:2021:AML

[GWQ21] Omid Gheibi, Danny Weyns, and Federico Quin. Applying machine learning in selfadaptive systems: a systematic literature review. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15 (3):9:1–9:37, September 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3469440.

Gounaris:2012:ELB

[GYP12] Anastasios Gounaris, Christos A. Yfoulis, and Norman W. Paton. Efficient load balancing in partitioned queries under random perturbations. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):5:1–5:??, April 2012. CODEN ????

ISSN 1556-4665 (print), 1556-4703 (electronic).

Gounaris:2008:CTA

[GYSD08] Anastasios Gounaris, Christos Yfoulis, Rizos Sakellariou, and Marios D. Dikaiakos. A control theoretical approach to self-optimizing block transfer in Web service grids. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3 (2):6:1–6:??, May 2008. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Gu:2025:CNM

[GZBC25] Tianlong Gu, Taihang Zhi, Xuguang Bao, and Liang Chang. Credible negotiation for multi-agent reinforcement learning in long-term coordination. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(1): 1:1–1:??, March 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3706110.

Habib:2013:ASW

[HAMR13] Irfan Habib, Ashiq Anjum, Richard Mcclatchey, and Omer Rana. Adapting scientific workflow structures using multi-objective optimization strategies. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(1): 4:1–4:??, April 2013. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Harnie:2014:PUA

[HBDD14] Dries Harnie, Elisa Gonzalez Boix, Theo D’Hondt, and Wolfgang De Meuter. Programming urban-area applications by exploiting public transportation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9 (2):8:1–8:??, July 2014. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Hassan:2021:DEM

[HBMA21] Sara Hassan, Rami Bahsoon, Leandro Minku, and Nour Ali. Dynamic evaluation of microservice granularity adaptation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 16(2):4:1– 4:??, June 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502724.

Higashino:2016:AGR

[HEC^{+}16] Wilson A. Higashino, Cédric Eichler, Miriam A. M. Capretz, Luiz F. Bittencourt, and Thierry Monteil. Attributed graph rewriting for complex event processing selfmanagement. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(3):19:1–19:??, September 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Herrmann:2010:SOS

[Her10]  Klaus Herrmann.  Self-

organized service placement in ambient intelligence environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(2): 6:1–6:??, May 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Hilaire:2008:AAA

[HKR08] Vincent Hilaire, Abder Koukam, and Sebastian Rodriguez. An adaptative agent architecture for holonic multi-agent systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(1): 2:1–2:??, March 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Hao:2013:ASO

[HL13] Jianye Hao and Ho-Fung Leung. Achieving socially optimal outcomes in multiagent systems with reinforcement social learning. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(3):15:1–15:??, September 2013. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Hu:2021:GTB

[HLLL21] Shuyue Hu, Chin-Wing Leung, Ho-Fung Leung, and Jiamou Liu. Gist trace-based learning: Efficient convention emergence from multilateral interactions.

ACM Transactions on Autonomous and Adaptive Systems (TAAS), 16(1):2:1–2:??, March 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3502199.

Hao:2015:MRS

[HLM15] Jianye Hao, Ho-Fung Leung, and Zhong Ming. Multiagent reinforcement social learning toward coordination in cooperative multiagent systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(4):20:1–20:??, January 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Hosseinmardi:2015:DSG

[HMF^{+}15] Homa Hosseinmardi, Akshay Mysore, Nicholas Farrow, Nikolaus Correll, and Richard Han. Distributed spatiotemporal gesture recognition in sensor arrays. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(3):17:1– 17:??, October 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Hang:2011:TSS

[HS11] Chung-Wei Hang and Munindar P. Singh. Trustworthy service selection and composition. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(1): 5:1–5:??, February 2011. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Hao:2018:ERE

[HSC^{+}18] Jianye Hao, Jun Sun, Guangyong Chen, Zan Wang, Chao Yu, and Zhong Ming. Efficient and robust emergence of norms through heuristic collective learning. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12 (4):23:1–23:??, January 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Herbert:2007:ACM

[HSL^{+}07] Douglas Herbert, Vinaitheerthan Sundaram, Yung-Hsiang Lu, Saurabh Bagchi, and Zhiyuan Li. Adaptive correctness monitoring for wireless sensor networks using hierarchical distributed run-time invariant checking. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(3): 8:1–8:??, September 2007. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Handte:2012:SSA

[HSM^{+}12] Marcus Handte, Gregor Schiele, Verena Matjuntke, Christian Becker, and Pedro José Marrón. 3PC: System support for adaptive peer-to-peer pervasive computing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):10:1–10:??, April 2012. CODEN ????

ISSN 1556-4665 (print), 1556-4703 (electronic).

Hezavehi:2020:USA

[HWA^{+}20] Sara M. Hezavehi, Danny Weyns, Paris Avgeriou, Radu Calinescu, Raffaela Mirandola, and Diego Perez-Palacin. Uncertainty in self-adaptive systems: a research community perspective. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(4):10:1–10:36, December 2020. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3487921.

Hofstadler:2017:ECN

[HWH^{+}17] Daniel Nicolas Hofstadler, Mostafa Wahby, Mary Katherine Heinrich, Heiko Hamann, Payam Zahadat, Phil Ayres, and Thomas Schmickl. Evolved control of natural plants: Crossing the reality gap for user-defined steering of growth and motion. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12 (3):15:1–15:??, October 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Iannucci:2018:MBR

[IA18] Stefano Iannucci and Sherif Abdelwahed. Model-based response planning strategies for autonomic intrusion protection. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(1):4:1– 4:??, May 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Ippoliti:2016:OAA

[IJDZ16] Dennis Ippoliti, Changjun Jiang, Zhijun Ding, and Xiaobo Zhou. Online adaptive anomaly detection for augmented network flows. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(3):17:1– 17:??, September 2016. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Jiang:2018:UCS

[JAJ^{+}18] Jiuchuan Jiang, Bo An, Yichuan Jiang, Donghui Lin, Zhan Bu, Jie Cao, and Zhifeng Hao. Understanding crowdsourcing systems from a multiagent perspective and approach. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(2):8:1– 8:??, November 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Jelasity:2011:SSM

[JB11] Márk Jelasity and Vilmos Bilicki. Scalable stealth mode P2P overlays of very small constant degree. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6 (4):27:1–27:??, October 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Jiang:2013:FAE

[JH13] Hao Jiang and Jason O.

Hallstrom. Fast, accurate event classification on resource-lean embedded sensors. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(2):11:1– 11:??, July 2013. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Johnson:2007:MHD

[JI07] Jeffrey H. Johnson and Pejman Iravani. The multilevel hypernetwork dynamics of complex systems of robot soccer agents. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2 (2):5:1–5:??, June 2007. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Jiang:2015:RTA

[JZL15] Yichuan Jiang, Yifeng Zhou, and Yunpeng Li. Reliable task allocation with load balancing in multiplex networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(1):3:1–3:??, March 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Klinglmayr:2012:SOS

[KB12] Johannes Klinglmayr and Christian Bettstetter. Selforganizing synchronization with inhibitory-coupled oscillators: Convergence and robustness. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(3):30:1– 30:??, September 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kraemer:2015:RLI

[KB15] Landon Kraemer and Bikramjit Banerjee. Reinforcement learning of informed initial policies for decentralized planning. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(4):18:1–18:??, January 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kalyvianaki:2014:ARP

[KCH14] Evangelia Kalyvianaki, Themistoklis Charalambous, and Steven Hand. Adaptive resource provisioning for virtualized servers using Kalman filters. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(2):10:1– 10:??, July 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kolan:2007:STD

[KD07] Prakash Kolan and Ram Dantu. Socio-technical defense against voice spamming. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(1):2:1–2:??, March 2007. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Khaluf:2016:MRS

[KD16] Yara Khaluf and Marco Dorigo. Modeling robot swarms using integrals of birth-death processes. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(2):8:1–8:??, July 2016. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Ko:2008:NCN

[KGJ08] Steven Y. Ko, Indranil Gupta, and Yookyung Jo. A new class of nature-inspired algorithms for self-adaptive peer-to-peer computing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(3):11:1– 11:??, August 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kota:2012:DAS

[KGJ12] Ramachandra Kota, Nicholas Gibbins, and Nicholas R. Jennings. Decentralized approaches for self-adaptation in agent organizations. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):1:1–1:??, April 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kinneer:2021:IRS

[KGL21] Cody Kinneer, David Garlan, and Claire Le Goues. Information reuse and stochastic search: Managing uncertainty in self-* systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15 (1):3:1–3:36, February 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3440119.

Khan:2006:AFE

[KIW06] Masood Mehmood Khan, Michael Ingleby, and Robert D. Ward. Automated facial expression classification and affect interpretation using infrared measurement of facial skin temperature variations. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1(1):91–113, September 2006. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kuze:2016:CLS

[KKK^{+}16] Naomi Kuze, Daichi Kominami, Kenji Kashima, Tomoaki Hashimoto, and Masayuki Murata. Controlling largescale self-organized networks with lightweight cost for fast adaptation to changing environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(2):9:1– 9:??, July 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kuze:2018:HOC

[KKK^{+}18a] Naomi Kuze, Daichi Kominami, Kenji Kashima, Tomoaki Hashimoto, and Masayuki Murata. Hierarchical optimal control method for controlling large-scale self-organizing networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(4):22:1– 22:??, January 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kuze:2018:SOC

[KKK^{+}18b] Naomi Kuze, Daichi Kominami, Kenji Kashima, Tomoaki Hashimoto, and Masayuki Murata. Self-organizing control mechanism based on collective decision-making for information uncertainty. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(1):7:1–7:??, May 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Khan:2016:DMF

[KLWS16] Muhammad Umer Khan, Shuai Li, Qixin Wang, and Zili Shao. Distributed multirobot formation and tracking control in cluttered environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(2):12:1– 12:??, July 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Koshutanski:2008:IAC

[KM08] Hristo Koshutanski and Fabio Massacci. Interactive access control for autonomic systems: From theory to implementation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3 (3):9:1–9:??, August 2008. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kurka:2019:KMS

[KPO19] David Burth Kurka, Jeremy Pitt, and Josiah Ober. Knowledge management for selforganised resource allocation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(1):1:1–1:??, September 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kar:2016:CRS

[KRM16] Pushpendu Kar, Arijit Roy, and Sudip Misra. Connectivity reestablishment in selforganizing sensor networks with dumb nodes. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10 (4):28:1–28:??, February 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Kantert:2016:CNE

[KTK^{+}16] Jan Kantert, Sven Tomforde, Melanie Kauder, Richard Scharrer, Sarah Edenhofer, Jörg Hähner, and Christian Müller-Schloer. Controlling negative emergent behavior by graph analysis at runtime. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(2):7:1–7:??, July 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Langford:2021:EDD

[LC21] Michael Austin Langford and Betty H. C. Cheng. Enki: a diversity-driven approach to test and train robust learningenabled systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15 (2):5:1–5:32, June 2021. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3460959.

Li:2019:TTE

[LCQB19] Wenjuan Li, Jian Cao, Shiyou Qian, and Rajkumar Buyya. TSLAM: a trust-enabled selflearning agent model for service matching in the cloud market. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(4):16:1– 16:??, July 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3317604.

Liao:2018:APM

[LCT^{+}18] Jianwei Liao, Zhigang Cai, François Trahay, Jun Zhou, and Guoqiang Xiao. Adaptive process migrations in coupled applications for exchanging data in local file cache. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(2):9:1–9:??, November 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Liu:2018:SAP

[LDC^{+}18] Xunyun Liu, Amir Vahid Dastjerdi, Rodrigo N. Calheiros, Chenhao Qu, and Rajkumar Buya. A stepwise autoprofiling method for performance optimization of streaming applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12 (4):24:1–24:??, January 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Labella:2006:DLG

[LDD06] Thomas H. Labella, Marco Dorigo, and Jean-Louis Deneubourg. Division of labor in a group of robots inspired by ants’ foraging behavior. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1 (1):4–25, September 2006. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Liu:2016:TCD

[LDL16] Linfeng Liu, Jingli Du, and Ye Liu. Topology control for diverse coverage in underwater wireless sensor networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(3):16:1– 16:??, September 2016. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Lewis:2015:SDA

[LEC^{+}15] Peter R. Lewis, Lukas Esterle, Arjun Chandra, Bernhard Rinner, Jim Torresen, and Xin Yao. Static, dynamic, and adaptive heterogeneity in distributed smart camera networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10 (2):8:1–8:??, June 2015. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Lee:2019:IDA

[LF19] Gil Jae Lee and José A. B.

Fortes. Improving dataanalytics performance via autonomic control of concurrency and resource units. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(3):13:1– 13:??, March 2019. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3309539.

Liang:2023:MDC

[LHAES23] Qianlin Liang, Walid A.

Hanafy, Ahmed Ali-Eldin, and Prashant Shenoy. Modeldriven cluster resource management for AI workloads in edge clouds. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18 (1):2:1–2:??, March 2023. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3582080.

Lesch:2023:SAO

[LHKK23] Veronika Lesch, Marius Hadry, Christian Krupitzer, and Samuel Kounev. Self-aware optimization of adaptation planning strategies. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(3):10:1–10:??, September 2023. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3568680.

Litoiu:2007:PAM

[Lit07]  Marin Litoiu.  A perfor-

mance analysis method for autonomic computing systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(1):3:1–3:??, March 2007. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

LeBlond:2012:CPB

[LLL12] Stevens Le Blond, Fabrice Le Fessant, and Erwan Le Merrer. Choosing partners based on availability in P2P networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(2):25:1– 25:??, July 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Legtchenko:2012:RCR

[LMSM12] Sergey Legtchenko, Sébastien Monnet, Pierre Sens, and Gilles Muller. RelaxDHT: a churn-resilient replication strategy for peer-to-peer distributed hash-tables. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(2):28:1–28:??, July 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Loureiro:2012:DOC

[LND12] Emerson Loureiro, Paddy Nixon, and Simon Dobson. Decentralized and optimal control of shared resource pools. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):14:1–14:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Li:2024:UGP

[LNS24] Jia Li, Shiva Nejati, and Mehrdad Sabetzadeh. Using genetic programming to build self-adaptivity into softwaredefined networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19 (1):2:1–2:??, March 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3616496.

Lemmon:2009:ISI

[LPZZ09] Michael Lemmon, Christian Poellabauer, Liqiang Zhang, and Xiaobo Zhou. Introduction to the special issue on selfadaptive and self-organizing wireless networking systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(3):15:1–15:??, July 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Leroux:2012:DOS

[LR12] Philippe Leroux and Sébastien Roy. Designing and optimizing swarming in a distributed base station network: Application to power control. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7 (2):24:1–24:??, July 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Lee:2009:IIA

[LS09] Chonho Lee and Junichi Suzuki. An immunologicallyinspired autonomic framework for self-organizing and evolvable network applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(4):22:1–22:??, November 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Li:2025:PWF

[LSZ25] Jiaye Li, Jiagang Song, and Shichao Zhang. Piecewise weighting function for collaborative filtering recommendation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(1): 6:1–6:??, March 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3708353.

Liu:2025:CBB

[LTS^{+}25] Bowen Liu, Hao Tian, Zhijie Shen, Yueyue Xu, and Wanchun Dou. A consortium blockchain-based edge task offloading method for connected autonomous vehicles. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(3):19:1– 19:27, September 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Locatelli:2007:ACU

[LV07] Marco P. Locatelli and Giuseppe Vizzari. Awareness in collaborative ubiquitous environments: The multilayered multi-agent situated system approach. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(4):13:1– 13:??, November 2007. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Loia:2010:ESI

[LV10] Vincenzo Loia and Athanasios V. Vasilakos. Editorial: Special issue on ambient intelligence. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5 (2):5:1–5:??, May 2010. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Lee:2015:DDC

[LVP15] Eun Kyung Lee, Hariharasudhan Viswanathan, and Dario Pompili. Distributed data-centric adaptive sampling for cyber-physical systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(4):21:1–21:??, January 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Li:2016:MSC

[LWQL16] Chao Li, Rui Wang, Depei Qian, and Tao Li. Managing server clusters on renewable energy mix. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11 (1):1:1–1:??, April 2016. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Luo:2014:MDA

[LXX^{+}14] Jiaqing Luo, Bin Xiao, Qingjun Xiao, Jiannong Cao, and Minyi Guo. Modeling and defending against adaptive BitTorrent worms in peer-to-peer networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(1): 5:1–5:??, March 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Lama:2013:APS

[LZ13] Palden Lama and Xiaobo Zhou. Autonomic provisioning with self-adaptive neural fuzzy control for percentile-based delay guarantee. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8 (2):9:1–9:??, July 2013. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Langford:2024:AMF

[LZC24] Michael Austin Langford, Sol Zilberman, and Betty Cheng. Anunnaki: a modular framework for developing trusted artificial intelligence. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(3):17:1– 17:??, September 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3649453.

Li:2024:GAS

[LZL^{+}24a] Jialong Li, Mingyue Zhang, Nianyu Li, Danny Weyns, Zhi Jin, and Kenji Tei. Generative AI for self-adaptive systems: State of the art and research roadmap. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(3):13:1– 13:??, September 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3686803.

Li:2024:GTS

[LZL^{+}24b] Nianyu Li, Mingyue Zhang, Jialong Li, Sridhar Adepu, Eunsuk Kang, and Zhi Jin. A game-theoretical self-adaptation framework for securing softwareintensive systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(2):12:1–12:??, June 2024.

CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3652949.

Mu:2018:SFE

[MAFS^{+}18] Ting-Yu Mu, Ala Al-Fuqaha, Khaled Shuaib, Farag M. Sallabi, and Junaid Qadir. SDN flow entry management using reinforcement learning. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(2):11:1– 11:??, November 2018. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mashayekhi:2022:PNE

[MALS22] Mehdi Mashayekhi, Nirav Ajmeri, George F. List, and Munindar P. Singh. Prosocial norm emergence in multi-agent systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 17(1–2):3:1– 3:??, June 2022. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3540202.

Mamei:2011:MPS

[Mam11] Marco Mamei. Macro programming a spatial computer with Bayesian networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(2):16:1–16:??, June 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mukhtar:2011:DUT

[MBB11] Hamid Mukhtar, Djamel Belaïd, and Guy Bernard. Dynamic user task composition based on user preferences. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(1):4:1–4:??, February 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Moreno:2018:FED

[MCGS18] Gabriel A. Moreno, Javier Cámara, David Garlan, and Bradley Schmerl. Flexible and efficient decision-making for proactive latency-aware selfadaptation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13 (1):3:1–3:??, May 2018. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mirandola:2025:ISS

[MCJ25] Raffaela Mirandola, Radu Calinescu, and Pooyan Jamshidi. Introduction to the SEAMS 2023 special issue. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(2):8:1–8:2, June 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mordacchini:2020:HCD

[MCPB20] Matteo Mordacchini, Marco Conti, Andrea Passarella, and Raffaele Bruno. Humancentric data dissemination in the IoP: Large-scale modeling and evaluation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14 (3):10:1–10:25, March 2020. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3366372.

Marinescu:2017:PBM

[MDC17] Andrei Marinescu, Ivana Dusparic, and Siobhán Clarke. Prediction-based multi-agent reinforcement learning in inherently non-stationary environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(2):9:1– 9:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mencagli:2016:GTA

[Men16] Gabriele Mencagli. A gametheoretic approach for elastic distributed data stream processing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(2):13:1– 13:??, July 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Maignan:2011:GGA

[MG11] Luidnel Maignan and Frédéric Gruau. Gabriel graphs in arbitrary metric space and their cellular automaton for many grids. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6 (2):12:1–12:??, June 2011. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Maggio:2012:CDM

[MHP^{+}12] Martina Maggio, Henry Hoffmann, Alessandro V. Papadopoulos, Jacopo Panerati, Marco D. Santambrogio, Anant Agarwal, and Alberto Leva. Comparison of decision-making strategies for self-optimization in autonomic computing systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(4):36:1– 36:??, December 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mellouk:2013:SDT

[MHZ13] Abdelhamid Mellouk, Said Hoceini, and Sherali Zeadally. A state-dependent time evolving multi-constraint routing algorithm. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(1): 6:1–6:??, April 2013. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mena:2006:SRS

[MIRG06] Eduardo Mena, Arantza Illarramendi, Jose A. Royo, and Alfredo GoñI. A software retrieval service based on adaptive knowledge-driven agents for wireless environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1(1):67–90, September 2006. CODEN ????

ISSN 1556-4665 (print), 1556-4703 (electronic).

Maliah:2021:CCP

[MKS21] Shlomi Maliah, Radimir Komarnitski, and Guy Shani. Computing contingent plan graphs using online planning. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 16(1):1:1–1:??, March 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3488903.

Metzger:2024:USE

[MLFP24] Andreas Metzger, Jan Laufer, Felix Feit, and Klaus Pohl. A user study on explainable online reinforcement learning for adaptive systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(3):15:1–15:??, September 2024. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3666005.

Morales:2015:OAS

[MLsRA^{+}15] Javier Morales, Maite Lópezsánchez, Juan A. Rodriguez-Aguilar, Wamberto Vasconcelos, and Michael Wooldridge. Online automated synthesis of compact normative systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(1):2:1–2:??, March 2015. CODEN ????

ISSN 1556-4665 (print), 1556-4703 (electronic).

Mali:2017:TMB

[MM17] Goutam Mali and Sudip Misra.

Topology management-based distributed camera actuation in wireless multimedia sensor networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(1):2:1– 2:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Misra:2021:SSR

[MOP21] Sudip Misra, Tamoghna Ojha, and Madhusoodhanan P. SecRET: Secure range-based localization with evidence theory for underwater sensor networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(1):2:1– 2:26, February 2021. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3431390.

Marin-Perianu:2010:AVC

[MPBMP^{+}10] Mihai Marin-Perianu, Stephan Bosch, Raluca Marin-Perianu, Hans Scholten, and Paul Havinga. Autonomous vehicle coordination with wireless sensor and actuator networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(4):13:1–13:??, November 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mordacchini:2015:CTC

[MPC^{+}15] Matteo Mordacchini, Andrea Passarella, Marco Conti, Stuart M. Allen, Martin J. Chorley, Gualtiero B. Colombo, Vlad Tanasescu, and Roger M. Whitaker. Crowdsourcing through cognitive opportunistic networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10 (2):13:1–13:??, June 2015. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Misra:2011:BFI

[MR11] Sudip Misra and Gopidi Rajesh. Bird flight-inspired routing protocol for mobile ad hoc networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6 (4):25:1–25:??, October 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mense:2012:ERE

[MS12] Mario Mense and Christian Schindelhauer. An erasureresilient encoding system for flexible reading and writing in storage networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7 (2):22:1–22:??, July 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Magalhaes:2015:SWS

[MS15] João Paulo Magalhães and Luis Moura Silva. SHõ WA: a self-healing framework for Web-based applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(1):4:1–4:??, March 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mansour:2009:IPC

[MSA09] Mohamed S. Mansour, Karsten Schwan, and Sameh Abdelaziz. Isolation points: Creating performance-robust enterprise systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4 (2):10:1–10:??, May 2009. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Masuzawa:2009:BTK

[MT09] Toshimitsu Masuzawa and Sébastien Tixeuil. On bootstrapping topology knowledge in anonymous networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(1):8:1–8:??, January 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mencagli:2014:CPC

[MVV14] Gabriele Mencagli, Marco Vanneschi, and Emanuele Vespa. A cooperative predictive control approach to improve the reconfiguration stability of adaptive distributed parallel applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(1): 2:1–2:??, March 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Mamei:2007:PPB

[MZ07] Marco Mamei and Franco Zambonelli. Pervasive pheromonebased interaction with RFID tags. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(2): 4:1–4:??, June 2007. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Nia:2025:EMV

[NCKA25] Mehran Alidoost Nia, Radu Calinescu, Mehdi Kargahi, and Alessandro Abate. Efficient model verification at runtime through adaptive dynamic approximation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20 (1):3:1–3:??, March 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3708560.

Nuseibeh:2018:EF

[Nus18] Bashar Nuseibeh. Editorial: The first. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(1):1:1– 1:??, May 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

OToole:2017:DDE

[ONC17] Eamonn O’Toole, Vivek Nallur, and Siobhán Clarke. Decentralised detection of emergence in complex adaptive systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(1):4:1– 4:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Purandare:2025:ESA

[PAS^{+}25] Salil Purandare, Md Nafee Al Islam, Urjoshi Sinha, Jane Cleland-Huang, and Myra B. Cohen. An evaluation of self-adaptive mechanisms for misconfigurations in small uncrewed aerial systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(2):9:1–9:35, June 2025. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Purkayastha:2013:CRA

[PB13] Punyaslok Purkayastha and John S. Baras. Convergence results for ant routing algorithms via stochastic approximation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(1): 3:1–3:??, April 2013. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Peleteiro:2014:FCT

[PBARA14] Ana Peleteiro, Juan C. Burguillo, Josep Ll. Arcos, and Juan A. Rodriguez-Aguilar. Fostering cooperation through dynamic coalition formation and partner switching. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(1):1:1–1:??, March 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Pitt:2014:DJS

[PBM14] Jeremy Pitt, Dídac Busquets, and Sam Macbeth. Distributive justice for self-organised common-pool resource management. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(3):14:1– 14:??, October 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Pfannemuller:2020:RIM

[PBW^{+}20] Martin Pfannemüller, Martin Breitbach, Markus Weckesser, Christian Becker, Bradley Schmerl, Andy Schürr, and Christian Krupitzer. REACT-ION: a model-based runtime environment for situationaware adaptations. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(4):12:1–12:29, December 2020. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3487919.

Porter:2025:ICF

[PCL^{+}25] Zoe Porter, Radu Calinescu, Ernest Lim, Victoria Hodge, Philippa Ryan, Simon Burton, Ibrahim Habli, Tom Lawton, John McDermid, John Molloy, Helen Monkhouse, Phillip Morgan, Paul Noordhof, Colin Paterson, Isobel Standen, and Jie Zou. INSYTE: a classification framework for traditional to agentic AI systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(3):15:1– 15:39, September 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Phithakkitnukoon:2011:BBA

[PDCE11] Santi Phithakkitnukoon, Ram Dantu, Rob Claxton, and Nathan Eagle. Behaviorbased adaptive call predictor. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(3):21:1–21:??, September 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Pitt:2015:SSR

[PH15] Jeremy Pitt and Tom Holvoet.

SASO 2013: Selected, revised, and extended best papers. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(2):7:1–7:??, June 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Pianini:2023:FAS

[PK23] Danilo Pianini and Vana Kalogeraki. Foreword: AC-SOS 2021 special issue. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(3):8:1–8:??, September 2023. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3612929.

Pang:2025:ASH

[PLQ^{+}25] Shengye Pang, Yi Li, Zhen Qin, Xinkui Zhao, Jintao Chen, Fan Wang, and Jianwei Yin. Adaptive scheduling of high-availability drone swarms for congestion alleviation in connected automated vehicles. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(3):18:1– 18:19, September 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Petta:2007:ISI

[POPM07] Paolo Petta, Andrea Omicini, Terry Payne, and Peter McBurney. Introduction to the special issue: The AgentLink III technical forums. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(4):12:1– 12:??, November 2007. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Poslad:2007:SPM

[Pos07]  Stefan Poslad.  Specifying

protocols for multi-agent systems interaction. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(4):15:1–15:??, November 2007. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Pournaras:2018:DCL

[PPA18] Evangelos Pournaras, Peter Pilgerstorfer, and Thomas Asikis. Decentralized collective learning for self-managed sharing economies. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(2):10:1–10:??, November 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Petruzzi:2017:ESC

[PPB17] Patricio E. Petruzzi, Jeremy Pitt, and Dídac Busquets. Electronic social capital for self-organising multi-agent systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(3):13:1– 13:??, October 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Pianini:2022:CAA

[PPCE22] Danilo Pianini, Federico Pettinari, Roberto Casadei, and Lukas Esterle. A collective adaptive approach to decentralised k-coverage in multirobot systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 17 (1–2):4:1–4:??, June 2022. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3547145.

Penserini:2007:HVD

[PPSM07] Loris Penserini, Anna Perini, Angelo Susi, and John Mylopoulos. High variability design for software agents: Extending tropos. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(4):16:1–16:??, November 2007. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Poola:2016:ERW

[PRB16] Deepak Poola, Kotagiri Ramamohanarao, and Rajkumar Buyya. Enhancing reliability of workflow execution using task replication and spot instances. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(4):30:1– 30:??, February 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Pei:2011:SOS

[PRJ11] Guanhong Pei, Binoy Ravindran, and E. Douglas Jensen. Self-organizing and self-reconfigurable event routing in ad hoc networks with causal dependency awareness. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(3):19:1– 19:??, September 2011. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Paiva:2015:ASS

[PRRR15] João Paiva, Pedro Ruivo, Paolo Romano, and Luís Rodrigues. AutoPlacer: Scalable self-tuning data placement in distributed key-value stores. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 9(4):19:1– 19:??, January 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Pitt:2012:ASE

[PSA12] Jeremy Pitt, Julia Schaumeier, and Alexander Artikis. Axiomatization of socio-economic principles for self-organizing institutions: Concepts, experiments and challenges. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(4):39:1–39:??, December 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Popescu:2012:FTD

[PSB^{+}12] Razvan Popescu, Athanasios Staikopoulos, Antonio Brogi, Peng Liu, and Siobhán Clarke. A formalized, taxonomydriven approach to crosslayer application adaptation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):7:1–7:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Puviani:2012:MFA

[PSFC12] Mariachiara Puviani, Giovanna Di Marzo Serugendo, Regina Frei, and Giacomo Cabri. A method fragments approach to methodologies for engineering self-organizing systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(3):33:1– 33:??, September 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Pani:2015:CSS

[PSPR15] Danilo Pani, Carlo Sau, Francesca Palumbo, and Luigi Raffo. Computing swarms for self-adaptiveness and selforganization in floating-point array processing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10 (3):16:1–16:??, October 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Paurobally:2007:FWS

[PTW07] Shamimabi Paurobally, Valentina Tamma, and Michael Wooldrdige. A framework for Web service negotiation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(4):14:1–14:??, November 2007. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Parashar:2011:E

[PZ11] Manish Parashar and Franco Zambonelli. Editorial. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(4):29:1–29:??, October 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Parashar:2013:E

[PZ13] Manish Parashar and Franco Zambonelli. Editorial. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(1):1:1–1:??, April 2013. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Parashar:2018:FE

[PZ18] Manish Parashar and Franco Zambonelli. Farewell editorial. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(4):17:1– 17:??, January 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Qi:2025:ISI

[QKSV25] Lianyong Qi, Burak Kantarci, Houbing Song, and Anna Maria Vegni. Introduction to the special issue on intelligent applications of Web 3.0 and Metaverse for connected autonomous vehicles. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(3):16:1– 16:4, September 2025. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Quiroz:2012:DED

[QPGS12] Andres Quiroz, Manish Parashar, Nathan Gnanasambandam, and Naveen Sharma. Design and evaluation of decentralized online clustering. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(3):34:1–34:??, September 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Rodriguez:2017:BDS

[RB17] Maria A. Rodriguez and Rajkumar Buyya. Budget-driven scheduling of scientific workflows in IaaS clouds with fine-grained billing periods. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(2):5:1–5:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Rzadca:2015:GTM

[RDKB15] Krzysztof Rzadca, Anwitaman Datta, Gunnar Kreitz, and Sonja Buchegger. Game-theoretic mechanisms to increase data availability in decentralized storage systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(3):14:1– 14:??, October 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Raza:2016:UIB

[RH16] Saleha Raza and Sajjad Haider. Using imitation to build collaborative agents. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(1):3:1–3:??, April 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Roy:2017:TCS

[RMKM17] Arijit Roy, Sudip Misra, Pushpendu Kar, and Ayan Mondal. Topology control for self-adaptation in wireless sensor networks with temporary connection impairment. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(4):21:1– 21:??, February 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Riganelli:2019:CIL

[RMM19] Oliviero Riganelli, Daniela Micucci, and Leonardo Mariani. Controlling interactions with libraries in Android apps through runtime enforcement. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(2):8:1–8:29, December 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3368087.

Rahimian:2015:DAL

[RPG^{+}15] Fatemeh Rahimian, Amir H.

Payberah, Sarunas Girdzijauskas, Mark Jelasity, and Seif Haridi. A distributed algorithm for large-scale graph partitioning. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(2):12:1– 12:??, June 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Rudolph:2019:MIA

[RTH19] Stefan Rudolph, Sven Tomforde, and Jörg Hähner. Mutual influence-aware runtime learning of self-adaptation behavior. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(1): 4:1–4:??, September 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3345319.

Rahman:2017:CAC

[RTN^{+}17] Muntasir Raihan Rahman, Lewis Tseng, Son Nguyen, Indranil Gupta, and Nitin Vaidya. Characterizing and adapting the consistency– latency tradeoff in distributed key–value stores. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11 (4):20:1–20:??, February 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Raza:2020:HFA

[RW20] Syed Ali Raza and Mary-Anne Williams. Human feedback as action assignment in interactive reinforcement learning. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(4):14:1– 14:24, September 2020. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3404197.

Ren:2007:RRS

[RYC^{+}07] Shangping Ren, Yue Yu, Nianen Chen, Jeffrey J.-P. Tsai, and Kevin Kwiat. The role of roles in supporting reconfigurability and fault localizations for open distributed and embedded systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(3):10:1– 10:??, September 2007. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Soundararajan:2006:RPB

[SA06] Gokul Soundararajan and Cristiana Amza. Reactive provisioning of backend databases in shared dynamic content server clusters. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1 (2):151–188, December 2006. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Sang:2012:SSF

[SA12] Lifeng Sang and Anish Arora.

A shared-secret free security infrastructure for wireless networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(2):23:1– 23:??, July 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Sarkadi:2024:SGH

[Sar24] Stefan Sarkadi. Self-governing hybrid societies and deception. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(2):9:1–9:??, June 2024. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3638549.

Sharma:2017:TAC

[SBMM17] Gokarna Sharma, Costas Busch, Supratik Mukhopadhyay, and Charles Malveaux. Tight analysis of a collisionless robot gathering algorithm. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(1):3:1–3:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Sabatucci:2019:SDW

[SC19] Luca Sabatucci and Massimo Cossentino. Supporting dynamic workflows with automatic extraction of goals from BPMN. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(2):7:1– 7:??, October 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Silva:2017:HLA

[SCC17] Fernando Silva, Luís Correia, and Anders Lyhne Christensen. Hyper-learning algorithms for online evolution of robot controllers. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12 (3):14:1–14:??, October 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Schmerl:2024:FSS

[SCM24] Bradley Schmerl, Javier Cámara, and Martina Maggio. Foreword: SEAMS 2022 special issue. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(1):1:1–1:??, March 2024. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643642.

Sunel:2024:FMB

[SÇP24] Saim Sunel, Erkin Çilden, and Faruk Polat. Faster MIL-based subgoal identification for reinforcement learning by tuning fewer hyperparameters. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(2): 10:1–10:??, June 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3643852.

Shams:2020:ABR

[SDOP20] Zohreh Shams, Marina De Vos, Nir Oren, and Julian Padget. Argumentation-based reasoning about plans, maintenance goals, and norms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14 (3):9:1–9:39, March 2020. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3364220.

Souissi:2009:UEC

[SDY09] Samia Souissi, Xavier Défago, and Masafumi Yamashita. Using eventually consistent compasses to gather memory-less mobile robots with limited visibility. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(1): 9:1–9:??, January 2009. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Serugendo:2006:I

[Ser06] Giovanna Di Marzo Serugendo. Introduction. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1 (1):1–3, September 2006. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Serbedzija:2012:RPS

[SF12] Nikola Serbedzija and Stephen Fairclough. Reflective pervasive systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7 (1):12:1–12:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Siqueira:2025:ESP

[SFdL25] Bento Rafael Siqueira, Fabiano Cutigi Ferrari, and Rogério de Lemos. An extended study of the performance of flexible controllers composed of micro-controllers. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(2):10:1– 10:41, June 2025. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Salvaneschi:2013:ALL

[SGP13] Guido Salvaneschi, Carlo Ghezzi, and Matteo Pradella. An analysis of language-level support for self-adaptive software. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(2): 7:1–7:??, July 2013. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Schuhmann:2013:ACD

[SHRB13] Stephan Schuhmann, Klaus Herrmann, Kurt Rothermel, and Yazan Boshmaf. Adaptive composition of distributed pervasive applications in heterogeneous environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(2):10:1– 10:??, July 2013. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Schmerl:2017:ISS

[SI17] Bradley Schmerl and Paola Inverardi. Introduction to the special section on best papers from SEAMS 2015. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(4):22:1– 22:??, February 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Semwal:2018:OMR

[SJN18] Tushar Semwal, Shashi Shekhar Jha, and Shivashankar B. Nair. On ordering multi-robot task executions within a cyber physical system. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12 (4):20:1–20:??, January 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Skandylas:2020:DTA

[SKA20] Charilaos Skandylas, Narges Khakpour, and Jesper Andersson. AT-DIFC +: Toward adaptive and trust-aware decentralized information flow control. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(4):13:1– 13:35, December 2020. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3487292.

Shen:2008:ABD

[SLJS08] Chien-Chung Shen, Ke Li, Chaiporn Jaikaeo, and Vinay Sridhara. Ant-based distributed constrained Steiner tree algorithm for jointly conserving energy and bounding delay in ad hoc multicast routing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(1):3:1–3:??, March 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Song:2025:MCW

[SLW25] Jian Song, Guanjun Liu, and Miaomiao Wang. Model checking of workflow nets with tables and constraints.

ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(2):14:1– 14:38, June 2025. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Sui:2015:AOD

[SMHP15] Zhiquan Sui, Matthew Malensek, Neil Harvey, and Shrideep Pallickara. Autonomous orchestration of distributed discrete event simulations in the presence of resource uncertainty. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10(3):18:1– 18:??, October 2015. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Sabuhi:2021:OPC

[SMK21] Mikael Sabuhi, Nima Mahmoudi, and Hamzeh Khazaei. Optimizing the performance of containerized cloud software systems using adaptive PID controllers. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(3):8:1– 8:27, September 2021. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3465630.

Schmeck:2010:ASO

[SMSÇ^{+}10] Hartmut Schmeck, Christian Müller-Schloer, Emre Çakar, Moez Mnif, and Urban Richter. Adaptivity and selforganization in organic computing systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(3):10:1–10:??, September 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Silva:2024:SAT

[SPB24] Samira Silva, Patrizio Pelliccione, and Antonia Bertolino. Self-adaptive testing in the field. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(1): 4:1–4:??, March 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3627163.

Shyu:2007:NID

[SQX^{+}07] Mei-Ling Shyu, Thiago Quirino, Zongxing Xie, Shu-Ching Chen, and Liwu Chang. Network intrusion detection through adaptive sub-eigenspace modeling in multiagent systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(3): 9:1–9:??, September 2007. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Such:2016:PPN

[SR16] Jose M. Such and Michael Rovatsos. Privacy policy negotiation in social media. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(1):4:1–4:??, April 2016. CODEN ????

ISSN 1556-4665 (print), 1556-4703 (electronic).

Saffre:2012:HST

[SS12] Fabrice Saffre and Aistis Simaitis. Host selection through collective decision. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):4:1–4:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Stratan:2012:XRS

[SSN^{+}12] Corina Stratan, Jan Sacha, Jeff Napper, Paolo Costa, and Guillaume Pierre. The XtreemOS resource selection service. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(4):37:1– 37:??, December 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Schmidt:2023:URS

[SSVB23] Jorge F. Schmidt, Udo Schilcher, Arke Vogell, and Christian Bettstetter. Using randomization in self-organized synchronization for wireless networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(3):9:1–9:??, September 2023. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3605553.

Salehie:2009:SAS

[ST09] Mazeiar Salehie and Ladan Tahvildari. Self-adaptive software: Landscape and research challenges. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(2):14:1– 14:??, May 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Schneider:2013:CSC

[ST13] Daniel Schneider and Mario Trapp. Conditional safety certification of open adaptive systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(2): 8:1–8:??, July 2013. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Shevtsov:2019:SCT

[SWM19] Stepan Shevtsov, Danny Weyns, and Martina Maggio. SimCA*: a control-theoretic approach to handle uncertainty in selfadaptive systems with guarantees. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(4):17:1– 17:??, July 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3328730.

Su:2019:IAA

[SZB19] Xing Su, Minjie Zhang, and Quan Bai. An innovative approach for ad hoc network establishment in disaster environments by the deployment of wireless mobile agents. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(4):19:1– 19:??, July 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3337795.

Su:2020:FLS

[SZY^{+}20] Zhaopin Su, Guofu Zhang, Feng Yue, Jindong He, Miqing Li, Bin Li, and Xin Yao. Finding the largest successful coalition under the strict goal preferences of agents. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(4):15:1–15:33, September 2020. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3412370.

Tuci:2006:CTS

[TGT^{+}06] Elio Tuci, Roderich Groß, Vito Trianni, Francesco Mondada, Michael Bonani, and Marco Dorigo. Cooperation through self-assembly in multirobot systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 1 (2):115–150, December 2006. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Terraneo:2025:CFF

[TKFL25] Federico Terraneo, Zaigham Khalid, William Fornaciari, and Alberto Leva. Calibrationfree feedforward temperature compensation for wireless clock synchronisation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(1):7:1–7:??, March 2025. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3716138.

Tacconi:2011:CES

[TMC^{+}11] David Tacconi, Daniele Miorandi, Iacopo Carreras, Francesco De Pellegrini, and Imrich Chlamtac. Cooperative evolution of services in ubiquitous computing environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(3):20:1–20:??, September 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Tunde-Onadele:2024:SSM

[TOLG^{+}24] Olufogorehan Tunde-Onadele, Yuhang Lin, Xiaohui Gu, Jingzhu He, and Hugo Latapie. Self-supervised machine learning framework for online container security attack detection. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(3):16:1– 16:??, September 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3665795.

Tsai:2007:ISI

[TS07] Jeffrey J. P. Tsai and Mukesh Singhal. Introduction: Special issue of the IEEE SUTC’06. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(3):7:1–7:??, September 2007. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Tong:2023:GGD

[TSLF23] Junbo Tong, Daming Shi, Yi Liu, and Wenhui Fan. GLDAP: Global Dynamic Action Persistence Adaptation for deep reinforcement learning. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(2):7:1–7:??, June 2023. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3590154.

Toosi:2016:AMC

[TVKB16] Adel Nadjaran Toosi, Kurt Vanmechelen, Farzad Khodadadi, and Rajkumar Buyya. An auction mechanism for cloud spot markets. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(1):2:1–2:??, April 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Tomforde:2020:ISI

[TWS20] Sven Tomforde, Timothy Wood, and Jan-Philipp Steghöfer. Introduction to the special issue with Selected Papers of The International Conference on Autonomic Computing and Self-Organizing Systems (AC-SOS) 2020. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(4):10e:1– 10e:2, December 2020. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3492340.

Urgaonkar:2008:ADP

[USC^{+}08] Bhuvan Urgaonkar, Prashant Shenoy, Abhishek Chandra, Pawan Goyal, and Timothy Wood. Agile dynamic provisioning of multi-tier Internet applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3 (1):1:1–1:??, March 2008. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Vu:2011:EUC

[VA11] Le-Hung Vu and Karl Aberer. Effective usage of computational trust models in rational environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6 (4):24:1–24:??, October 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Viroli:2011:SCP

[VCMZ11] Mirko Viroli, Matteo Casadei, Sara Montagna, and Franco Zambonelli. Spatial coordination of pervasive services through chemical-inspired tuple spaces. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(2):14:1– 14:??, June 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Viroli:2016:SSR

[VDK16] Mirko Viroli, Ada Diaconescu, and Nagarajan Kandasamy. SASO 2014: Selected, revised, and extended best papers. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(2):5:1–5:??, July 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Vogel:2014:MDE

[VG14] Thomas Vogel and Holger Giese. Model-driven engineering of self-adaptive software with EUREMA. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8 (4):18:1–18:??, January 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Vrancx:2015:RLA

[VGR^{+}15] Peter Vrancx, Pasquale Gurzi, Abdel Rodriguez, Kris Steenhaut, and Ann Nowé. A reinforcement learning approach for interdomain routing with link prices. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10 (1):5:1–5:??, March 2015. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

VonKistowski:2017:MEL

[VHK^{+}17] Jóakim Von Kistowski, Nikolas Herbst, Samuel Kounev, Henning Groenda, Christian Stier, and Sebastian Lehrig. Modeling and extracting load intensity profiles. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11 (4):23:1–23:??, February 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Venkatasubramanian:2014:CAP

[VMG14] Krishna K. Venkatasubramanian, Tridib Mukherjee, and Sandeep K. S. Gupta. CAAC — an adaptive and proactive access control approach for emergencies in smart infrastructures. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8 (4):20:1–20:??, January 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Vasilakos:2009:ESI

[VP09] Athanasios V. Vasilakos and Witold Pedrycz. Editorial to the special issue. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(4):20:1–20:??, November 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Villatoro:2013:RCE

[VSMS13] Daniel Villatoro, Jordi Sabater-Mir, and Sandip Sen. Robust convention emergence in social networks through selfreinforcing structures dissolution. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(1):

2:1–2:??, April 2013. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wareham:2019:DRT

[War19] Todd Wareham. Designing robot teams for distributed construction, repair, and maintenance. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(1):2:1– 2:??, September 2019. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wurtz:2010:ESI

[WBSI10] Rolf P. Würtz, Kirstie L. Bellman, Hartmut Schmeck, and Christian Igel. Editorial: Special issue on organic computing. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(3):9:1–9:??, September 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wang:2009:SOF

[WCD^{+}09] Yu Wang, Lijuan Cao, Teresa A.

Dahlberg, Fan Li, and Xinghua Shi. Self-organizing faulttolerant topology control in large-scale three-dimensional wireless networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(3):19:1–19:??, July 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wang:2017:IRL

[WCW^{+}17] Hongbign Wang, Xin Chen, Qin Wu, Qi Yu, Xingguo Hu, Zibin Zheng, and Athman Bouguettaya. Integrating reinforcement learning with multi-agent techniques for adaptive service composition. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(2):8:1–8:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

White:2011:SSD

[WDTS11] Jules White, Brian Dougherty, Chris Thompson, and Douglas C. Schmidt. ScatterD: Spatial deployment optimization with hybrid heuristic/evolutionary algorithms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(3):18:1–18:??, September 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Weyns:2023:SAI

[WGA^{+}23] Danny Weyns, Ilias Gerostathopoulos, Nadeem Abbas, Jesper Andersson, Stefan Biffl, Premek Brada, Tomas Bures, Amleto Di Salle, Matthias Galster, Patricia Lago, Grace Lewis, Marin Litoiu, Angelika Musil, Juergen Musil, Panos Patros, and Patrizio Pelliccione. Selfadaptation in industry: a survey. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 18(2):5:1–5:??, June 2023. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3589227.

Weyns:2022:DLE

[WGQV22] Danny Weyns, Omid Gheibi, Federico Quin, and Jeroen Van Der Donckt. Deep learning for effective and efficient reduction of large adaptation spaces in self-adaptive systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 17(1–2):1:1– 1:??, June 2022. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3530192.

Weyns:2025:AVB

[WHA^{+}25] Danny Weyns, Sara M.

Hezavehi, Paris Avgeriou, Radu Calinescu, Raffaela Mirandola, and Diego Perez-Palacin. An architectural viewpoint for benefit-cost-risk-aware decision-making in selfadaptive systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20 (1):2:1–2:??, March 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3705612.

Weyns:2010:MOM

[WHH10a] Danny Weyns, Robrecht Haesevoets, and Alexander Helleboogh. The MACODO organization model for contextdriven dynamic agent organizations. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(4):16:1– 16:??, November 2010. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Weyns:2010:MMC

[WHH^{+}10b] Danny Weyns, Robrecht Haesevoets, Alexander Helleboogh, Tom Holvoet, and Wouter Joosen. The MA-CODO middleware for contextdriven dynamic agent organizations. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(1): 3:1–3:??, February 2010. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wang:2024:REA

[WHY24] Ziming Wang, Changwu Huang, and Xin Yao. A roadmap of explainable artificial intelligence: Explain to whom, when, what and how? ACM Transactions on Autonomous and Adaptive Systems (TAAS), 19(4):20:1– 20:??, December 2024. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3702004.

Wu:2025:SCL

[WLC^{+}25] Xiaotong Wu, Yuwen Liu, Xiaoxiao Chi, Rong Jiang, Xiaokang Zhou, Wajid Rafique, and Maqbool Khan. Secure collaborative learning for self-adaptive systems on connected autonomous vehicles. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(3):20:1– 20:30, September 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Weyns:2012:FUR

[WMA12] Danny Weyns, Sam Malek, and Jesper Andersson. FORMS: Unifying reference model for formal specification of distributed self-adaptive systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(1):8:1–8:??, April 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Watanabe:2007:RFP

[WNET07] Kenichi Watanabe, Yoshio Nakajima, Tomoya Enokido, and Makoto Takizawa. Ranking factors in peer-to-peer overlay networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 2(3):11:1–11:??, September 2007. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wang:2012:HPK

[WNV12a] Yufeng Wang, Akihiro Nakao, and Athanasios V. Vasilakos. Heterogeneity playing key role: Modeling and analyzing the dynamics of incentive mechanisms in autonomous networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(3):31:1– 31:??, September 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wang:2012:MCS

[WNV12b] Yufeng Wang, Akihiro Nakao, and Athanasios V. Vasilakos. On modeling of coevolution of strategies and structure in autonomous overlay networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(2):17:1–17:??, July 2012. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wang:2010:EBT

[WS10] Yonghong Wang and Munindar P. Singh. Evidence-based trust: a mathematical model geared for multiagent systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(4):14:1–14:??, November 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wang:2018:ECM

[WUK^{+}18] Cheng Wang, Bhuvan Urgaonkar, George Kesidis, Aayush Gupta, Lydia Y. Chen, and Robert Birke. Effective capacity modulation as an explicit control knob for public cloud profitability. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13(1):2:1–2:??, May 2018. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wareham:2018:VAO

[WV18] Todd Wareham and Andrew Vardy. Viable algorithmic options for designing reactive robot swarms. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 13 (1):5:1–5:??, May 2018. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wang:2017:SMC

[WVT^{+}17] Yang Wang, Bharadwaj Veeravalli, Chen-Khong Tham, Shuibing He, and Chengzhong Xu. On service migrations in the cloud for mobile accesses: a distributed approach. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 12(2):6:1–6:??, May 2017. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Wen:2025:SSA

[WXG^{+}25] Linfeng Wen, Minxian Xu, Sukhpal Singh Gill, Muhammad Hilman, Satish Narayana Srirama, Kejiang Ye, and Chengzhong Xu. StatuScale: Status-aware and elastic scaling strategy for microservice applications. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20 (1):4:1–4:??, March 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3686253.

Wang:2010:PSO

[WXZ10] Yu-Xuan Wang, Qiao-Liang Xiang, and Zhen-Dong Zhao. Particle swarm optimizer with adaptive tabu and mutation: a unified framework for efficient mutation operators. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 5(1):1:1–1:??, February 2010. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Xu:2025:MAR

[XGB^{+}25] Xiaolong Xu, Linjie Gu, Muhammad Bilal, Maqbool Khan, Yiping Wen, Guoqiang Liu, and Yuan Yuan. Multiagent reinforcement learning based edge content caching for connected autonomous vehicles in IoV. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(3):17:1– 17:26, September 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Xu:2012:PPB

[XLX12] Shouhuai Xu, Wenlian Lu, and Li Xu. Push- and pull-based epidemic spreading in networks: Thresholds and deeper insights. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7(3):32:1– 32:??, September 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Xu:2014:AED

[XLXZ14] Shouhuai Xu, Wenlian Lu, Li Xu, and Zhenxin Zhan. Adaptive epidemic dynamics in networks: Thresholds and control. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(4):19:1– 19:??, January 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Xiong:2011:APA

[XVYH11] Naixue Xiong, Athanasios V.

Vasilakos, Laurence T. Yang, and Ekram Hossain. An adaptive and predictive approach for autonomic multirate multicast networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(3):22:1–22:??, September 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Xu:2009:MLD

[XWN09] Bo Xu, Ouri Wolfson, and Channah Naiman. Machine learning in disruption-tolerant MANETs. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(4):23:1– 23:??, November 2009. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Xiao:2011:PIC

[XZL11] Yang Xiao, Yanping Zhang, and Xiannuan Liang. Primateinspired communication methods for mobile and static sensors and RFID tags. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 6(4):26:1–26:??, October 2011. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Yuan:2014:SSS

[YEM14] Eric Yuan, Naeem Esfahani, and Sam Malek. A systematic survey of self-protecting software systems. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8 (4):17:1–17:??, January 2014. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Yen:2016:DSS

[YHT16] Li-Hsing Yen, Jean-Yao Huang, and Volker Turau. Designing self-stabilizing systems using game theory. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 11(3):18:1–18:??, September 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Yu:2008:AAT

[YTW08] Zhenwei Yu, Jeffrey J. P. Tsai, and Thomas Weigert. An adaptive automatically tuning intrusion detection system. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 3(3):10:1–10:??, August 2008. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Yang:2025:WEM

[YZS^{+}25] Yihong Yang, Zhangbing Zhou, Lei Shu, Feng Zhou, Walid Gaaloul, and Arif Ali Khan. Web 3.0-enabled microservice re-scheduling for heterogenous resources cooptimization in metaverseintegrated edge networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(3):22:1– 22:26, September 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Zhang:2012:CDT

[ZCS12] Kai Zhang, Emmanuel G.

Collins, Jr., and Dongqing Shi. Centralized and distributed task allocation in multi-robot teams via a stochastic clustering auction. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7 (2):21:1–21:??, July 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Zhang:2013:PMO

[ZCVL13] Zhuoyao Zhang, Ludmila Cherkasova, Abhishek Verma, and Boon Thau Loo. Performance modeling and optimization of deadline-driven Pig programs. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 8(3):14:1– 14:??, September 2013. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Zhang:2020:UVP

[ZHSP20] Ruiwen Zhang, Tom Holvoet, Bifeng Song, and Yang Pei. UAVs vs. Pirates: an anticipatory swarm monitoring method using an adaptive pheromone map. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 14(4):13:1– 13:31, September 2020. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

URL https://dl.acm.org/
doi/10.1145/3380782.

Zhu:2021:VSF

[ZLHC21] Changxi Zhu, Ho-Fung Leung, Shuyue Hu, and Yi Cai. A Q-values sharing framework for multi-agent reinforcement learning under budget constraint. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 15(2):4:1– 4:28, June 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447268.

Zaker:2022:FVS

[ZLS22] Farzin Zaker, Marin Litoiu, and Mark Shtern. Formally verified scalable look ahead planning for cloud resource management. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 17(3–4):6:1–6:??, December 2022. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3555315.

Zou:2025:CPF

[ZLW^{+}25] Guobing Zou, Shiyi Lin, Shaogang Wu, Shengxiang Hu, Song Yang, Yanglan Gan, Bofeng Zhang, and Yixin Chen. Combining personalized federated hypernetworks and shared residual learning for distributed QoS prediction. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 20(3):23:1– 23:25, September 2025. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Zudaire:2021:AMA

[ZNU21] Sebastián A. Zudaire, Leandro Nahabedian, and Sebastián Uchitel. Assured mission adaptation of UAVs. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 16(3–4):7:1–7:??, December 2021. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3513091.

Zambonelli:2012:ISS

[ZP12] Franco Zambonelli and Ben Paechter. Introduction to the special section on pervasive adaptation. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 7 (1):9:1–9:??, April 2012. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Zhang:2009:MAA

[ZS09] Zonghua Zhang and Hong Shen. M-AID: An adaptive middleware built upon anomaly detectors for intrusion detection and rational response. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(4):24:1– 24:??, November 2009. CO-DEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Zhang:2009:CSD

[ZSA09] Hongwei Zhang, Lifeng Sang, and Anish Arora. On the convergence and stability of datadriven link estimation and routing in sensor networks. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 4(3):18:1–18:??, July 2009. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).

Zoghi:2016:DAA

[ZSLG16] Parisa Zoghi, Mark Shtern, Marin Litoiu, and Hamoun Ghanbari. Designing adaptive applications deployed on cloud environments. ACM Transactions on Autonomous and Adaptive Systems (TAAS), 10 (4):25:1–25:??, February 2016. CODEN ???? ISSN 1556-4665 (print), 1556-4703 (electronic).